### v1

# Measuring Informativeness Gap of (Mis)Calibrated Predictors

Abstract
In many applications, decision-makers must choose between multiple predictive models that may all be miscalibrated. Which model (i.e., predictor) is more “useful” in downstream decision tasks? To answer this, our first contribution introduces the notion of the informativeness gap between any two predictors, defined as the maximum normalized payoff advantage one predictor offers over the other across all decision-making tasks. Our framework strictly generalizes several existing notions: it subsumes U-Calibration (Kleinberg et al., 2023) and Calibration Decision Loss (Hu and Wu, 2024), which compare a miscalibrated predictor to its calibrated counterpart, and it recovers Blackwell informativeness (Blackwell, 1951, 1953) as a special case when both predictors are perfectly calibrated. Our second contribution is a dual characterization of the informativeness gap, which gives rise to a natural informativeness measure that can be viewed as a relaxed variant of the earth mover’s distance (EMD) between two prediction distributions. We show that this measure satisfies natural desiderata: it is complete and sound, and it can be estimated sample-efficiently in the prediction-only access setting. Along the way, we also obtain novel combinatorial structural results when applying this measure to perfectly calibrated predictors.
1 Introduction
Over the last decade, the machine learning predictors have grown remarkably powerful, especially with the rapid advancements in large-scale models such as large language models (LLMs). These predictors have demonstrated strong performance across a wide range of domains, providing high-quality predictions that are increasingly used by downstream decision-makers to inform their decisions. However, there are usually two key challenges that often hinder decision-makers from fully leveraging these predictions: (1) the underlying mechanisms used to generate predictions are frequently proprietary and opaque to external users; (2) due to limitations in training data or computational constraints in the training process, these predictors may exhibit biases and fail to accurately reflect the empirical frequencies of outcomes. To mitigate these challenges, one natural solution is to ensure that the predictions are calibrated.
A calibrated predictor regulates that predicted probabilities align with the true (conditional) probability of the outcome (Dawid, 1982; Foster and Vohra, 1998). For example, predictions of “80% likelihood” materialize approximately 80% outcome realizations of the time. It is well established that agents who naively best respond to perfectly calibrated predictions incur no regret (Foster and Vohra, 1998; Foster and Hart, 2021). With this desired property, a variety of calibration error metrics have been proposed to quantify how much a predictor deviates from perfect calibration – such as the Expected Calibration Error (ECE) (Foster and Vohra, 1997), the smooth calibration error (Foster and Hart, 2018), and the distance to calibration (Blasiok et al., 2023). Remarkably, some “decision-theoretic” measures are proposed to directly quantify the decision-makers’ regret when she best-responds to the predictor’s forecasts. Kleinberg et al. (2023) introduced “U-Calibration” (UCal), a measure that captures the maximum payoff loss incurred by an agent who naively best responds to a miscalibrated predictor, compared to responding to a best calibrated predictor that provides a fixed prediction. Similarly, Hu and Wu (2024) proposed the “Calibration Decision Loss” (CDL) which quantifies the maximum payoff gap between naively responding to a miscalibrated predictor and responding to the true empirical frequencies of the outcomes associated with each prediction.
Both UCal and CDL therefore assess how far a predictor falls short of its own calibrated counterpart – they tell us “how much better this predictor would be if it were perfectly calibrated.” In practice, however, the decision-maker is often handed two (or more) distinct predictors, each trained separately, each potentially miscalibrated in its own way, and must decide which one to rely on. Because neither predictor is necessarily the calibrated version of the other, existing metrics do not directly answer the questions like “Which of these two (possibly miscalibrated) predictors will yield the higher payoff? Is a perfectly calibrated predictor always more useful than a miscalibrated predictor?” The answers to these questions are far from obvious. To illustrate this subtlety, we begin with a simple weather-forecasting example.
Example 1.1.
Suppose the long‐run probability of rain is 50%. We compare following predictors:
  • The predictor $\nu$ always forecasts a $50\%$ chance of rain.
  • The predictor $\mu_{1}$ forecasts four possible predictions $\{0\%,49\%,51\%,100\%\}$, with a conditional prediction distribution constructed as
    0% $\mathbf{49\%}$ $\mathbf{51\%}$ 100%
    When it rains 00 $0.051$ $0.049$ 0.99
    When it does not rain 0.99 $0.049$ $510.0$
  • The predictor $\mu_{2}$ forecasts two possible predictions $\{1\%,99\%\}$, with a conditional prediction distribution constructed as
    $\mathbf{1\%}$ $\mathbf{99\%}$
    When it rains 00 $1$
    When it does not rain $1$ 00
In this example, we can see that although the predictor $\nu$ is perfectly calibrated but it contains no information about the realized outcome beyond the base rate (empirical rainy frequency). In contrast, although predictor $\mu_{1}$ miscalibrates slightly on the two middle predictions – it says $49\%$ (resp. $51\%$) but the true rain rate is $51\%$ (resp. $49\%$) – but is otherwise nearly perfectly accurate: it predicts 100% almost surely when it rains, and 0% almost surely when it does not. Because of this, any decision‐maker who acts on predictor $\mu_{1}$’s predictions can make choices that more closely reflect the actual outcome, compared to relying solely on the base rate. In fact, and perhaps unsurprisingly, one can show that for every decision problem – no matter the payoff structure – the expected payoff under miscalibrated predictor $\mu_{1}$ is never lower (and often strictly higher) than under perfectly calibrated predictor $\nu$. This example demonstrates that a miscalibrated but more informative predictor can dominate a perfectly calibrated yet uninformative one.
However, this dominance is not guaranteed in general. If we slightly alter predictor $\mu_{1}$ to be predictor $\mu_{2}$ such that it miscalibrates on the two extreme predictions – i.e., predicting $99\%$ when it actually rains, and predicting $1\%$ when it actually does not rain – then its advantage no longer holds. [^1] In fact, one can show that the ECEs of predictors $\mu_{1}$ and $\mu_{2}$ have the same magnitude. In this case, in some decision problems, the predictor $\mu_{2}$ leads to strictly lower expected payoff than predictor $\nu$, despite still being more “informative” in some sense.
The above example highlights that not all miscalibrations are equal in terms of their impact on decision-making. This raises a natural question: Which predictor is more useful for decision-making problems? Notably, this question is not fully answered even when both predictors are perfectly calibrated. In particular, by viewing a predictor as an information structure, the Blackwell informativeness order offers a partial answer to this question (Blackwell, 1951, 1953). Intuitively, the Blackwell informativeness and its induced Blackwell order captures whether one perfectly calibrated predictor is always more useful than another in every decision problem, thereby inducing a partial order over the space of perfectly calibrated predictors. However, not every pair of perfectly calibrated predictors is comparable under the Blackwell order, let alone pairs of possibly miscalibrated predictors. This work aims to study the following questions:
Can we compare any two (possibly miscalibrated) predictors based on how “useful” they are to the decision-making problems?
If they are not equally useful, can we quantify their gap, and is there a natural measure that satisfies common sense desiderata for characterizing this difference?
1.1 Main Results
In this paper, we provide principled answers to both of the motivating questions. In line with most prior work in the calibration literature, we focus on predictors for binary outcomes.
Informativeness gap.  As our first conceptual contribution, we introduce and study a new notion called the informativeness gap, denoted by $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$. Given any two (possibly miscalibrated) predictors $\mu$ and $\nu$, the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ quantifies the maximum payoff advantage that predictor $\mu$ offers over predictor $\nu$ across all normalized decision-making tasks. Here, the normalized decision-making tasks refer to those in which, for any action, the decision-maker’s payoff difference over different outcomes are bounded by one (see the formal definition in Definition 2.2). [^2] Alternatively, one could consider a multiplicative informativeness gap, defined as the maximum ratio between the payoffs under two predictors across all decision problems. However, we focus on and study the additive version, as it aligns more naturally with the decision-theoretic calibration literature, which primarily emphasizes additive regret.
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]\triangleq\underset{\begin{subarray}{c}\textrm{payoff-normalized}\\ \textrm{decision problem}\end{subarray}}{\sup}\textsc{Payoff}\!\left[{\mu}\right]-\textsc{Payoff}\!\left[{\nu}\right]~{},$
where $\textsc{Payoff}\!\left[{\mu}\right]$ denotes the expected utility in a particular decision problem obtained by the decision-maker who naively best-responds to the predictions produced by predictor. This definition provides an operational analogue to Blackwell informativeness, which interprets the relative usefulness of two predictors through the decision-maker’s maximum payoff difference on all possible decision problems.
Our proposed informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ also subsumes several existing decision-theoretic notions of calibration. For instance, U-Calibration corresponds to the special case where predictor $\mu$ always outputs the true base rate, while CDL arises when $\mu$ outputs the true conditional empirical frequencies induced by predictor $\nu$. When both $\mu$ and $\nu$ are perfectly calibrated, the Blackwell order offers a partial characterization of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$. In particular, we have $\textsc{InfoGap}\!\left[{\mu,\nu}\right]=0$ if and only if $\nu$ is more Blackwell informative than $\mu$. In other words, the Blackwell order provides an ordinal, partial comparison–indicating which predictor is more informative–without quantifying how much more informative one is than the other. However, the Blackwell framework does not extend to miscalibrated predictors, nor does it yield a cardinal measure of informativeness. The main contribution of this paper is to go beyond these limitations by extending the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ to any pair of predictors–including miscalibrated ones–while providing a meaningful cardinal measure of their relative usefulness.
Characterizing $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ between perfectly calibrated predictors.  We begin by analyzing the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ in the special but theoretically fundamental case where both predictors $\mu$ and $\nu$ are perfectly calibrated. Our main result provides a characterization of this gap, and interestingly, reveals that it closely resembles a relaxed variant of the well-known earth mover’s distance (EMD), also known as Wasserstein distance, between probability distributions. Motivated by this connection, we introduce the quantity $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$, defined over probability distributions. We then illustrate how to interpret it as an informativeness measure that quantifies how much more useful predictor $\mu$ is compared to predictor $\nu$ in decision-making tasks.
Definition 3.2 and Theorem 3.1 (Informal). For any two distributions $f_{1},f_{2}$, define the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ as
$\textsc{REMD}\!\left[{f_{1},f_{2}}\right]\triangleq\inf\nolimits_{\pi\in\Pi(f_{1},f_{2})}\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p$
and $\Pi(f_{1},f_{2})$ denotes the set of all couplings between two distributions. For any two perfectly calibrated predictors $\mu,\nu$, their informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$
where $f_{\mu},f_{\nu}\in\Delta([0,1])$ denote the distributions over predictions generated by the two predictors $\mu,\nu$, respectively.
While the informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ is defined over predictors, our relaxed earth mover’s distance $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ is defined over distributions. As we discuss below, $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ admits interesting structural characterizations and may be of independent interest—even beyond the context of calibration or forecasting. [^3] Throughout the paper, we use $f_{\mu}$ and $f_{\nu}$ to denote the prediction distributions associated with predictors $\mu$ and $\nu$, respectively, and $f_{1}$ and $f_{2}$ to denote general distributions.
By definition, we always have $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]\geq 0$ for all distributions $f_{1},f_{2}$. When $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]=0$ for prediction distributions $f_{\mu}$ and $f_{\nu}$, our results guarantee that, for all decision problems, the decision-maker obtains weakly higher utility under the perfectly calibrated predictor $\nu$ than under $\mu$. In this case, $\nu$ is more Blackwell informative than $\mu$. Conversely, when $\mu$ is more Blackwell informative than $\nu$, the value $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$ quantifies the extent to which $\mu$ is more informative than $\nu$, in terms of the maximal utility gap achievable across all payoff-normalized decision problems.
Since $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$ can be viewed as a relaxation of the classical earth mover’s distance (EMD), its value is always weakly smaller than that of the EMD. Given their structural similarity, one might wonder whether the EMD could serve as a proxy for the informativeness gap. Unfortunately, this is not the case: it is not difficult to construct examples where the multiplicative gap between the $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$ and the $\textsc{EMD}\!\left[{f_{\mu},f_{\nu}}\right]$ of two prediction distributions $f_{\mu},f_{\nu}$ is arbitrarily large (see Example 3.3).
To establish Theorem 3.1, we first formulate $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ as an infinite-dimensional linear program (LP), and then show that it admits a strong duality with the measure $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$. Along the way, we also derive alternative and insightful representations of $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$, which may be of independent interest.
Proposition 3.5 and Proposition 3.7 (Informal). For two distributions $f_{1},f_{2}$ with the same mean and supports contained in $[0,1]$, $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ admits the following two representations:
  • SCDFs representation: $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]=2\cdot\max_{t\in[0,1]}S_{1}(t)-S_{2}(t)$ where $S_{1},S_{2}$ denote the super-cumulative distribution functions (SCDFs) of distributions $f_{1},f_{2}$, respectively – that is, the integrals of their cumulative distribution functions. Also see Figure 1.
  • Lattice representation: $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]=\textsc{REMD}\!\left[{f_{1}\vee f_{2},f_{2}}\right]$ where $f_{1}\vee f_{2}$ denotes the join of distributions $f_{1}$ and $f_{2}$ under the Blackwell partial order. Also see Figure 2.
While the above representations are stated for $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ over distributions, by invoking the equivalence between $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ and $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ (Theorem 3.1), we also obtain the analogous results for the informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ over predictors. Specifically, the SCDF representation shows that $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ is exactly twice the maximum gap between the SCDFs of the two prediction distributions. This offers a simple yet precise way to quantify the informativeness gap when the predictors are perfectly calibrated. It is known that the space of all perfectly calibrated predictors, together with the Blackwell order, forms a lattice structure (Kertz and Rösler, 2000; Müller and Scarsini, 2006). Our lattice representation provides a qualitative interpretation of this structure: $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ can be viewed as the informativeness gap between predictor $\nu$ and the least upper bound (join) of $\mu$ and $\nu$ in the lattice.
Characterizing $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ between miscalibrated predictors.  We now turn to the setting where both predictors may be miscalibrated. To capture the informativeness gap in the presence of miscalibration, it is essential to incorporate the true frequencies of outcomes conditional on predictions. Specifically, for any prediction $p\sim\mu$, we let $\kappa_{\mu}(p)\in[0,1]$ denote the true outcome frequency given the prediction. Our main result introduces a strict extension of the measure $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$. This generalized measure, denoted by $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{1},\kappa_{2}}\right]$, takes as input two distributions and two corresponding outcome functions.
Definition 4.1 and Theorem 4.1 (Informal). For any two distributions $f_{1},f_{2}\in\Delta([0,1])$ and two functions $\kappa_{1},\kappa_{2}:[0,1]\rightarrow[0,1]$, define informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{1},\kappa_{2}}\right]$ as
$\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa,\kappa_{2}}\right]\triangleq\inf\limits_{\pi\in\overline{\Pi}(f_{1},f_{2})}~{}$ $\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)(p-q)\,{\mathrm{d}}q+(\kappa_{1}(p)-p)f_{1}(p)-(\kappa_{2}(p)-p)f_{2}(p)\right|\,{\mathrm{d}}p~{},$
where the set $\overline{\Pi}(\mu,\nu)$, referred to as the flow set is a strict superset of the coupling set $\Pi(f_{1},f_{2})$, which imposes “flow conservation” constraint over $\pi\in\Delta([0,1]\times[0,1])$.
For any two (possibly miscalibrated) predictors $\mu,\nu$, their informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$
where $f_{\mu},f_{\nu}$ (resp. $\kappa_{\mu},\kappa_{\nu}$) are the prediction distributions (resp. true outcome frequencies) of predictors $\mu,\nu$, respectively.
To understand how $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ generalizes $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$, notice that when both predictors $\mu,\nu$ are perfectly calibrated, we have $\kappa_{\mu}(p)=\kappa_{\nu}(p)=p$ for all $p\in[0,1]$. In this case, the objective in $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ reduces exactly to that of $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$. Moreover, the flow set $\overline{\Pi}(f_{\mu},f_{\nu})$ (see Section 4 for the formal definition) generalizes the standard coupling set $\Pi(f_{\mu},f_{\nu})$, and thus it represents a relaxed constraint.
Interestingly, by Theorem 3.1 and Theorem 4.1, we can see that when both predictors are perfectly calibrated, the two measures equal to each other: $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]=\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$. In other words, optimizing over the relaxed flow set is also sufficient to characterize the informativeness gap in the perfectly calibrated setting.
Given the close connection between the flow set $\overline{\Pi}(f_{\mu},f_{\nu})$ and the coupling set $\Pi(f_{\mu},f_{\nu})$, one might wonder whether simply optimizing over $\Pi(f_{\mu},f_{\nu})$ in $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ would suffice – or at least yield a good approximation – to the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$. However, this is not the case: there exist predictors $\mu,\nu$ for which the gap between optimizing over $\Pi(f_{\mu},f_{\nu})$ in $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ and $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ is arbitrarily large (see Proposition 4.3). To prove this result, we also extend the SCDF representation in Proposition 3.5 to miscalibrated predictors in Lemma 4.4.
Desiderata of our informativeness gap and informativeness measure.  Błasiok et al. (2023) proposed the following desiderata for an ideal calibration measure that should satisfy: (1) the prediction-only accessibility: the measure can be evaluated by only querying the pair of prediction and its associated realized outcome; (2) the consistency: the measure is robust completeness (correct predictions have low error) and robust soundness (incorrect predictions have high error); and (3) the sample efficiency: the measure can be computed within accuracy $\varepsilon$ in time $\text{poly}(\nicefrac{{1}}{{\varepsilon}})$ with using $\text{poly}(\nicefrac{{1}}{{\varepsilon}})$ samples of pair (prediction, realized outcome).
Our proposed informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ can be served as a tool for quantifying the informativeness gap between (possibly miscalibrated) predictors. By Theorem 4.1, $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ is both complete and sound (in fact, it satisfies these criteria exactly), and thus it is consistent. By our definition, our proposed informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ and its equivalence representation, i.e., informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$, are also prediction-only accessible: notice that for any predictor $\mu$, its true conditional frequency $\kappa_{\mu}(p)$ can be computed using only the sample pair (prediction, realized outcome). Lastly, we also present sample complexity bounds for estimating the measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ and thus the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$. The algorithm for this estimation utilizes our structural characterizations for $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$, and has a sample complexity of $\text{poly}(\nicefrac{{1}}{{\varepsilon}})$ (see Theorem 5.1).
Finally, another desiderata that is usually considered in the machine learning is the continuity of a measure: the measure should be continuous w.r.t. the prediction distribution. This raises a natural question: does there exist an informativeness measure that satisfy the above three desiderata and this additional continuity property simultaneously? Perhaps not surprisingly, the answer is no. We present a general impossibility result showing that no informativeness measure can satisfy consistency (i.e., completeness and soundness) and continuity simultaneously.
Proposition 4.9 (Informal). For any informativeness measure over predictors, at least one of the following must fail: it is complete, it is sound, and it is continuous.
Finally, we note that the informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$, informativeness measures $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$, and $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ also connect to other well-known calibration measures like ECE, distance to calibration, which we elaborate on these connections in Section 4.2.
1.2 Related Work
Our work contributes to a growing line of research on calibration in decision-making settings.
Informativeness of predictors.  The goal of this paper is to quantify and compare the informativeness gap of any given pair of predictors. To our best knowledge, the study most closely related to ours is Garg et al. (2019) who study how to compare the information content of multiple predictors. However, their main focus is understand how a calibrated predictor’s information content shapes fairness outcomes, while our focus, by contrast, is a general characterization of the informativeness gap itself, independent of any particular fairness criterion. Beyond the machine learning community, viewing the predictors as some information sources, the theoretical economics and statistics literature has long been concerned with comparing the value of information sources (see, e.g., Blackwell, 1951, 1953; Lehmann, 1988; Kertz and Rösler, 2000; Müller and Scarsini, 2006; Mu et al., 2021; Brooks et al., 2022, 2024). These works study Blackwell experiments, which in our terminology correspond to perfectly calibrated predictors. To the best of our knowledge, this paper is the first to extend the notion of informativeness to predictors that may be miscalibrated.
Decision aspects of (miscalibrated) predictors.  Beginning with Dawid (1982); Foster and Vohra (1998), it is well established that agents who best respond to calibrated predictions (i.e., those with low ECE) achieve diminishing swap regret. Recent works have extended this foundation by introducing new decision-theoretic calibration measures that offer finer-grained guarantees on the regret incurred by downstream agents.
In particular, Kleinberg et al. (2023) introduce “U-Calibration” – a measure bounded by a small constant times the ECE, and show that sublinear U-Calibration is both necessary and sufficient for ensuring sublinear regret for all downstream agents. Roth and Shi (2024); Hu and Wu (2024) consider a similar setting but focusing on swap regret. In particular, Roth and Shi (2024) show that by ensuring unbiasedness relative to a carefully chosen collection of events, one can achieve diminishing swap regret for arbitrary downstream agents with an improved rate compared to using calibrated predictions. Hu and Wu (2024) then introduce “Calibration Decision Loss (CDL)”, defined as the maximum improvement (over all decision tasks) in decision payoff obtainable by recalibrating the predictions. The show that CDL is upper bounded by twice the ECE and that a vanishing CDL guarantees vanishing payoff loss from miscalibration across all decision tasks, which removes the regret dependence on the number of actions appeared in the results of Roth and Shi (2024). Qiao and Zhao (2025) propose a new calibration measure called “subsampled step calibration”, which has both decision-theoretic implications and a truthfulness property. These works mostly focus the decision loss of a given predictor compared to a reference calibrated predictor, while our focus is to compare the decision loss of any given two (possibly miscalibrated) predictors.
Other conceptually related works include Noarov et al. (2023); Haghtalab et al. (2023); Jain and Perchet (2024); Feng and Tang (2025); Tang et al. (2025) who focus on single decision task, and Camara et al. (2020); Collina et al. (2024) who explore repeated principal-agent interactions in a prior-free setting. Zhao et al. (2021) introduce “decision calibration” that requires the predicted distribution and true distribution to be indistinguishable to a set of downstream decision-makers. Gopalan et al. (2022) develop an omniprediction framework which aims to design a single predictor that ensures every downstream decision-maker’s loss is no worse than some benchmark family of functions (Gopalan et al., 2023; Garg et al., 2024; Okoroafor et al., 2025).
Calibration error measures.  In addition to above calibration measures that are tailored to decision-making environments, there are also various measures proposed to capture the calibration errors. For example, the aforementioned ECE metric is broadly accepted in the forecasting community (Dawid, 1982; Foster and Vohra, 1998; Gneiting et al., 2007; Ranjan and Gneiting, 2010), within theoretical computer science (see, e.g., Qiao and Valiant, 2021; Dagan et al., 2024; Casacuberta et al., 2024), empirical machine learning community (see, e.g., Gneiting and Katzfuss, 2014; Rahaman et al., 2021). Beyond the standard ECE, other calibration measures have been proposed to capture different desired aspects, e.g., multi-class calibration (Gopalan et al., 2024), distance to calibration (Błasiok et al., 2023; Qiao and Zheng, 2024; Arunachaleswaran et al., 2025).
2 Preliminaries
In this work, we are interested in the predictors for the the binary classification task. We explain their connection to decision making and their informativeness in Sections 2.1 and 2.2, respectively.
2.1 Predictions for Downstream Decision Making
Suppose there is a binary classification task, where the goal is to predict an unknown binary outcome $y\in\{0,1\}\triangleq\mathcal{Y}$. A predictor $\mu$ for this task encodes a distribution $f_{\mu}\in\Delta([0,1])$ over predictions $p\in[0,1]$. A realized prediction $p$ can be interpreted as the “predicted” probability that the outcome $y$ equals one. With a slight abuse of notation, we use $f_{\mu}(p)$ to denote the probability density (or mass) function (PDF) of the predictor $\mu$’s prediction distribution. We also denote by $F_{\mu}$ as the corresponding cumulative distribution function (CDF). [^4] Previous work such as Błasiok et al. (2023) models a predictor $\mu$ as a stochastic mapping from a given feature vector $x$ to a predicted probability $p$ of the outcome $y$. In contrast, we do not explicitly define the feature space or introduce $x$ in our model. The goal of this work is to compare the informativeness of different predictors, which may originate from machine learning models trained on different datasets with potentially incomparable or heterogeneous feature configurations. Moreover, treating the predictor as a black box—without requiring access to its input features—is often desirable in practice. This setting is known as the prediction-only access model (Błasiok et al., 2023), and our formulation is naturally aligned with it. See Section 5 for more discussions.
We say a predictor $\mu$ is perfectly calibrated if, conditional on its realized prediction $p$, the true probability of the outcome being one is exactly equal to $p$. Formally,
$\forall p\in\mathsf{supp}(f_{\mu}):\qquad\text{Pr}\!\left[{y=1\,\mid\,p}\right]=p~{},$ (perfectly calibrated)
where $\mathsf{supp}(f_{\mu})$ denotes the support of the distribution $f_{\mu}$, i.e., the set of realized predictions that occur with positive probability.
In most practical machine learning scenarios, perfect calibration is rarely achievable. Predictors are often miscalibrated, meaning their predicted probabilities do not match the true conditional probabilities. For a given predictor $\mu$, we define the true expected outcome [^5] Since the outcome $y$ is binary, the probability that it equals one is equal to its expectation. function conditional on the prediction $p$ as
$\kappa_{\mu}(p)\triangleq\text{Pr}\!\left[{y=1\mid p}\right].$
As a sanity check, a predictor $\mu$ is perfectly calibrated if and only if $\kappa_{\mu}(p)=p$ for all $p\in\mathsf{supp}(f_{\mu})$.
Decision making with prediction.  Suppose there is a downstream decision maker (henceforth, the agent) who uses predictions to inform his decisions. The agent’s decision problem is characterized by an action set $\mathcal{A}$ and a utility function $u:\mathcal{A}\times\mathcal{Y}\rightarrow\mathbb{R}$, which maps each action $a\in\mathcal{A}$ and realized outcome $y\in\mathcal{Y}$ to a utility received.
The agent uses the realized prediction $p$ generated by the predictor to inform his decision. Following the recent literature (Haghtalab et al., 2023; Garg et al., 2024; Kleinberg et al., 2023; Hu and Wu, 2024; Roth and Shi, 2024; Qiao and Zhao, 2025), we define the agent’s best response $a^{*}(p)\in\mathcal{A}$ when the agent trusts the provided prediction $p$. Specifically, this is the action that maximizes the agent’s expected utility assuming the outcome $y$ equals one with probability $p$:
$a^{*}(p)\in\operatorname{argmax}\nolimits_{a\in\mathcal{A}}~{}\mathbb{E}_{y\sim\mathsf{Bern}(p)}\!\left[{u(a,y)}\right]=\operatorname{argmax}\nolimits_{a\in\mathcal{A}}~{}p\cdot u(a,1)+(1-p)\cdot u(a,0)$ (1)
where $\mathsf{Bern}(p)$ denotes the Bernoulli distribution with mean $p$.
Let $\tilde{u}(p,y)\triangleq u(a^{*}(p),y)$ denote the utility the agent receives upon taking the best response to prediction $p$ when the true outcome is $y\in\mathcal{Y}$. Given any predictor $\mu$, the agent’s expected utility (also referred to as payoff) when always best-responding to realized predictions is given by:
$\textsc{Payoff}\!\left[{\mu}\right]\triangleq\mathbb{E}_{y}\!\left[{\mathbb{E}_{p\sim f_{\mu}}\!\left[{\tilde{u}(p,y)}\right]}\right]~{}.$
We further define the agent’s interim utility when best responding to a prediction $p$, while the true outcome is drawn from the Bernoulli distribution $\mathsf{Bern}(q)$:
$U(q;p)\triangleq q\cdot\tilde{u}(p,1)+(1-q)\cdot\tilde{u}(p,0)~{}.$
Using this definition, the agent’s expected utility under predictor $\mu$ can be equivalently expressed in terms of the interim utility function:
Lemma 2.1.
Given any predictor $\mu$, the agent’s expected utility $\textsc{Payoff}\!\left[{\mu}\right]$ when he best responds to every prediction $p\sim f_{\mu}$ can be formulated as
$\textsc{Payoff}\!\left[{\mu}\right]$ $=\displaystyle\int_{0}^{1}f_{\mu}(p)\cdot U(\kappa_{\mu}(p);p)\,{\mathrm{d}}p~{}.$
Consequently, if predictor $\mu$ is perfectly calibrated, the agent’s expected utility $\textsc{Payoff}\!\left[{\mu}\right]$ is
$\textsc{Payoff}\!\left[{\mu}\right]$ $=\int_{0}^{1}f_{\mu}(p)\cdot U(p)\,{\mathrm{d}}p~{},$
where $U(p)\triangleq U(p;p)$ denotes the univariate form of the agent’s interim utility function.
2.2 Blackwell Order and Informativeness Gap between Predictors
Suppose we are given two predictors, $\mu$ and $\nu$, along with their respective true expected outcome functions, $\kappa_{\mu}$ and $\kappa_{\nu}$. As we discussed in the introduction, natural and important question arises: can we compare these predictors based on how informative they are?
A partial answer to this question is provided by the notion of informativeness introduced by Blackwell (1951, 1953), which applies when the predictors are perfectly calibrated. In particular, a perfectly calibrated predictor can be viewed as an information structure—that is, a distribution over posterior beliefs about the binary outcome $y$ induced by some signaling scheme. [^6] A signaling scheme is a stochastic mapping from the realized outcome $y$ to a signal from a given signal space. Upon observing a realized signal, a posterior belief about the outcome can be formed via Bayes’ rule. Since the outcome is binary, the posterior belief is fully captured by the probability that the outcome is one. Thus, the distribution of posterior beliefs corresponds to a distribution over the probabilities of the outcome being one, with randomness induced by both the outcome and the signaling scheme. Such a distribution of posterior beliefs is mathematically equivalent to a perfectly calibrated predictor.
We now provide the formal definition of Blackwell informativeness:
Definition 2.1 (Blackwell informativeness, adopted from Blackwell, 1951, 1953).
Fix any two perfectly calibrated predictors $\mu,\nu$. We say predictor $\mu$ is Blackwell more informative than predictor $\nu$ (equivalently, $\mu$ Blackwell-dominates $\nu$), if for all decision problems $(\mathcal{A},u)$ of the agent,
$\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu}\right]\geq\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\nu}\right]~{},\quad\text{for all }(\mathcal{A},u)~{},$
where $\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu}\right]$ and $\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\nu}\right]$ are the agent’s expected utilities under predictors $\mu,\nu$ in decision problem $(\mathcal{A},u)$, respectively.
The concept of Blackwell informativeness is intuitive. In particular, this definition captures how perfectly calibrated predictors (information structures) enhance the decision maker’s utility for all decision problems. A perfectly calibrated predictor is Blackwell more informative than another if the induced utility in the former is strictly greater than the induced utility in the latter.
The definition of Blackwell informativeness above introduces a partial order over the space of perfectly calibrated predictors, known as the Blackwell order, which has been extensively studied in the literature. Both the Blackwell informativeness and the Blackwell order exhibit well-known properties. Below, we highlight two of these properties:
Lemma 2.2 (Blackwell, 1951, 1953).
For any two perfectly calibrated predictors $\mu$ and $\nu$, predictor $\mu$ Blackwell-dominates predictor $\nu$ if and only if for all $t\in[0,1]$,
$S_{\mu}(t)\geq S_{\nu}(t)~{}.$
Here $S_{\mu}$ and $S_{\nu}$ are the super-cumulative distribution functions (SCDFs) defined as
$S_{\mu}(t)\triangleq\displaystyle\int_{0}^{t}F_{\mu}(p)\,{\mathrm{d}}p\;\;\mbox{and}\;\;S_{\nu}(t)\triangleq\displaystyle\int_{0}^{t}F_{\nu}(t)\,{\mathrm{d}}p~{},$
where $F_{\mu}$ and $F_{\nu}$ are the CDFs of prediction distribution $f_{\mu}$ and $f_{\nu}$, respectively.
Lemma 2.3 (Kertz and Rösler, 2000; Müller and Scarsini, 2006).
The space of all perfectly calibrated predictors, along with the Blackwell order over this space, forms a lattice.
We remark on two key limitations of the Blackwell informativeness relation. First, it is only defined over the space of perfectly calibrated predictors, which can be viewed as information structures. Second, even within this restricted space, the Blackwell relation induces only a partial order, providing an ordinal comparison without quantifying how much more informative one predictor is than another.
The main focus of this work is to move beyond these limitations and develop a framework to compare the informativeness of any two predictors, including those that may be miscalibrated, in a cardinal manner. To this end, we introduce the following notion of the informativeness gap.
Definition 2.2 (Informativeness gap).
Given any two (possibly miscalibrated) predictors $\mu$ and $\nu$, the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ of predictor $\nu$ relative to predictor $\mu$ is defined as
$\begin{array}[]{rlll}\textsc{InfoGap}\!\left[{\mu,\nu}\right]\triangleq\sup\nolimits_{(\mathcal{A},u)\in\mathcal{U}}~{}\displaystyle\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu}\right]-\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\nu}\right]~{},\end{array}$ (3)
where $\mathcal{U}$ denotes the class of all decision problems $(\mathcal{A},u)$ such that for every action $a\in\mathcal{A}$, the utility difference is bounded:
$|u(a,1)-u(a,0)|\leq 1~{}.$
Notably, the bounded utility difference in Definition 2.2 serves primarily as a normalization. [^7] Previous work such as Kleinberg et al. (2023); Hu and Wu (2024) consider decision problems $(\mathcal{A},u)$ under the assumption that utilities are bounded in $[0,1]$, i.e., $u(a,0),u(a,1)\in[0,1]$ for all $a\in\mathcal{A}$. In contrast, our bounded utility difference condition in Definition 2.2 can be viewed as a relaxation of this assumption. Despite this restriction, the class of decision problems remains rich enough for the informativeness gap to fully capture the original notion of Blackwell informativeness, which considers all utility functions. [^8] Besides Proposition 2.4, we also establish connections between the informativeness gap and other properties (Lemmas 2.2 and 2.3) of Blackwell informativeness. See Propositions 3.5 and 3.7. Its proof can be found in Section A.1.
Proposition 2.4.
For any two perfectly calibrated predictors $\mu$ and $\nu$, predictor $\mu$ Blackwell-dominates predictor $\nu$ if and only if the informativeness gap of predictor $\mu$ relative to predictor $\nu$ is zero, i.e., $\textsc{InfoGap}\!\left[{\nu,\mu}\right]=0$.
We conclude this section with the following technical lemma regarding the agent’s interim utility function. Its proof can be found in Section A.2.
Lemma 2.5.
For every decision problem $(\mathcal{A},u)$, its induced univariate interim utility function $U(p)$ (see definition in Lemma 2.1) is 1-Lipschitz and convex in $p\in[0,1]$ if and only if for every action $a\in\mathcal{A}$, the utility difference is bounded, i.e., $|u(a,1)-u(a,0)|\leq 1$.
3 The Informativeness of Perfectly Calibrated Predictors
In this section, we focus on the case where both $\mu$ and $\nu$ are perfectly calibrated predictors. Our main result in this section is a dual characterization of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$. To state our main results, we first revisit a natural $\ell_{1}$-distance measure between two distributions.
Definition 3.1 (EMD between distributions).
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$, the earth mover’s distance $\textsc{EMD}\!\left[{f_{1},f_{2}}\right]$ is defined as follows:
$\textsc{EMD}\!\left[{f_{1},f_{2}}\right]\triangleq\inf\nolimits_{\pi\in\Pi(f_{1},f_{2})}\int_{0}^{1}\int_{0}^{1}\pi(p,q)\cdot\left|p-q\right|\,{\mathrm{d}}q{\mathrm{d}}p~{},$
where
$\Pi(f_{1},f_{2})\triangleq\left\{\pi\in\Delta\left([0,1]\times[0,1]\right):\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}q=f_{1}(p),\int_{0}^{1}\pi(q,p)\,{\mathrm{d}}q=f_{2}(p),\forall p\in[0,1]\right\}$
is referred to as the coupling set, and it denotes the set of all couplings for distributions $f_{1},f_{2}$.
We note that earth mover’s distance $\textsc{EMD}\!\left[{\mu,\nu}\right]$ is also known as the Wasserstein distance and has appeared in previous works (Błasiok et al., 2023; Qiao and Zheng, 2024), where it is referred to as the “lower distance” between predictors’ prediction distributions. The core of our characterization is a variant of the EMD, which we define as follows:
Definition 3.2 (Relaxed EMD between distributions).
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$, the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ is defined as follows:
$\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ $\triangleq\inf\nolimits_{\pi\in\Pi(f_{1},f_{2})}\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p~{}.$ (4)
We remark that unlike $\textsc{EMD}\!\left[{\cdot,\cdot}\right]$ which can be used as a metric in the space of all distributions, our proposed $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ is asymmetric as generally $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]\neq\textsc{REMD}\!\left[{f_{2},f_{1}}\right]$ and cannot be used as a metric. We are now ready to state our main result, which establishes the equivalence between the informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ and our proposed $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$.
Theorem 3.1.
For any two perfectly calibrated predictors $\mu,\nu$, their informativeness gap satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]~{}.$
where $f_{\mu}$ and $f_{\nu}$ are the prediction distributions of predictors $\mu$ and $\nu$, respectively.
By the triangle inequality, we always have $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]\leq\textsc{EMD}\!\left[{f_{1},f_{2}}\right]$. One tempting thought is that whether one can approximate $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ using the classic $\textsc{EMD}\!\left[{f_{\mu},f_{\nu}}\right]$ (e.g., whether they are polynomially related with each other). Perhaps not surprisingly, one can show that the multiplicative gap between $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ and $\textsc{EMD}\!\left[{f_{\mu},f_{\nu}}\right]$ can be arbitrarily large.
Example 3.3.
Consider perfectly calibrated predictor $\mu$ outputs deterministic prediction $p=0.5$, and perfectly calibrated predictor $\nu$ outputs a prediction $p\sim U[0,1]$ uniformly at random. Under this example, predictor $\nu$ Blackwell-dominates $\mu$, thus $\textsc{InfoGap}\!\left[{\mu,\nu}\right]=0$, while $\textsc{EMD}\!\left[{f_{\mu},f_{\nu}}\right]=0.25$.
In the remainder of this section, we first provide the formal proof of Theorem 3.1 in Section 3.1. We then establish additional structural characterization results of $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ in Section 3.2.
3.1 Proof of Theorem 3.1
The proof of Theorem 3.1 consists of three main steps:
  • $\bullet$
    In step 1 (see Lemma 3.2), we reformulate the informativeness gap as an optimization program over all the 1-Lipschitz convex interim utility functions.
  • $\bullet$
    In step 2 (see Lemma 3.3), we further cast program 7 to a new program (see program 11) that is more tractable to establish the equivalence to $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$.
  • $\bullet$
    In step 3 (see Lemma 3.4), we use strong duality to prove the equivalence between the program 11 and $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$.
Lemma 3.2.
For any two perfectly calibrated predictors $\mu,\nu$, their informativeness gap satisfies
[…]
$\begin{array}[]{rlll}\sup\limits_{U:[0,1]\rightarrow\mathbb{R}}{}&\displaystyle\int_{0}^{1}U(p)\cdot\left(f_{1}(p)-f_{2}(p)\right)~{}{\mathrm{d}}p&\vspace{1mm}\text{s.t.}\\ &\displaystyle U\text{ is $1$-Lipschitz convex},&\vspace{1mm}\end{array}$ (7)
Proof.
Invoking Lemmas 2.1 and 2.5 and the definition of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ completes the proof of Lemma 3.2. ∎
Lemma 3.3.
For any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$, the optimal objective value of program $\textsc{P}_{1}[f_{1},f_{2}]$ satisfies
[…]
$\begin{array}[]{rlll}\sup\limits_{\alpha_{1},\alpha_{2},\gamma:[0,1]\rightarrow\mathbb{R}}{}&\displaystyle\int_{0}^{1}\left(\alpha_{1}(p)+\gamma(p)\cdot p\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)f_{2}(q)\,{\mathrm{d}}q&\text{s.t.}&\vspace{1mm}\\ &\displaystyle\alpha_{1}(p)+\alpha_{2}(q)\leq-\gamma(p)\cdot q,&p,q\in[0,1]\vspace{1mm}\\ &\displaystyle\gamma(p)\in[-1,1],&p\in[0,1]\vspace{1mm}\end{array}$ (11)
Proof.
[…]
$\forall q\in[0,1]:\qquad V(q)\triangleq\inf\nolimits_{p\in[0,1]}-\gamma(p)\cdot q-\alpha_{1}(p)$
Notice that function $V$ is the highest lower concave envelope over affine functions $\{l_{p}(\cdot)\}_{p\in[0,1]}$ where each $l_{p}(q)\triangleq-\gamma(p)\cdot q-\alpha_{1}(p)$. This implies that function $V$ must be 1-Lipschitz concave, since the constraint that $|\gamma(\cdot)|\leq 1$ in program 11.
With the above definition of auxiliary function $V$, we can upper bound the objective function in program 11 as follows:
$\int_{0}^{1}\left(\alpha_{1}(p)+\gamma(p)\cdot p\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)\cdot f_{2}(q)\,{\mathrm{d}}q$ $\leq\int_{0}^{1}\left(\alpha_{1}(p)+\gamma(p)\cdot p\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}V(q)\cdot f_{2}(q)\,{\mathrm{d}}q$
$\leq\int_{0}^{1}-V(p)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}V(q)\cdot f_{2}(q)\,{\mathrm{d}}q$
$=\int_{0}^{1}-V(p)\cdot\left(f_{1}(p)-f_{2}(p)\right)\,{\mathrm{d}}p$
[…]
$U(p)\geq U(q)+\gamma(q)\cdot(p-q)~{}.$ (12)
Indeed, such slope function $\gamma$ can be chosen by picking any sub-gradient $\gamma(p)\in\partial U(p)$ (which exists a.e. for convex function $U$). The $1$-Lipschitzness of function $U$ ensures the second constraint that $|\gamma(\cdot)|\leq 1$ in program 11 is satisfied. Given $U$ and a choice of $\gamma$, we construct $\alpha_{1}$ and $\alpha_{2}$ as follows:
$\forall p\in[0,1]:\qquad\alpha_{1}(p)\triangleq U(p)-\gamma(p)\cdot p\;\;\mbox{and}\;\;\alpha_{2}(p)\triangleq-U(p)$
By construction, inequality (12) can be rewritten as
$-\alpha_{2}(p)\geq\alpha_{1}(q)+\gamma(q)\cdot p\;\;\Longleftrightarrow\;\;\alpha_{1}(q)+\alpha(p)\leq-\gamma(q)\cdot p$
and thus the first constraint in program 11 is satisfied. Thus, $\alpha_{1},\alpha_{2},\gamma$ is a feasible solution to program 11. Moreover, the objective value of program 11 under solution $\alpha_{1},\alpha_{2},\gamma$ satisfies that
$\int_{0}^{1}\left(\alpha_{1}(p)+\gamma(p)\cdot p\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)\cdot f_{2}(q)\,{\mathrm{d}}q$
[…]
Putting the two pieces together, we complete the proof of Lemma 3.3. ∎
Lemma 3.4.
For any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$, the the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ satisfies
[…]
Proof.
We prove the lemma statement by the strong duality. Recall the identity $\left|x\right|=\max_{\gamma\in[-1,1]}\gamma\cdot x$. Thus, given any feasible coupling $\pi\in\Pi(f_{1},f_{2})$, we can reformulate the objective function in program (4) in the definition of $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ as follows:
$\int_{0}^{1}\left|\int_{0}^{1}(p-q)\cdot\pi(p,q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p$ $=\int_{0}^{1}\left|p\cdot f_{1}(p)-\int_{0}^{1}q\cdot\pi(p,q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p$
$=\max_{\gamma:[0,1]\rightarrow[-1,1]}\int_{0}^{1}\gamma(p)\cdot\left(p\cdot f_{1}(p)-\int_{0}^{1}q\cdot\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p$
$=\max_{\gamma:[0,1]\rightarrow[-1,1]}\int_{0}^{1}\gamma(p)\cdot p\cdot f_{1}(p)\,{\mathrm{d}}p-\int_{0}^{1}\gamma(p)\cdot\left(\int_{0}^{1}q\cdot\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p~{}.$
We next consider Lagrange multipliers $\alpha_{1},\alpha_{2}:[0,1]\rightarrow\mathbb{R}$, and $\beta:[0,1]\times[0,1]\rightarrow\mathbb{R}_{+}$ for program (4):
$\mathcal{L}$ $=\int_{0}^{1}\gamma(p)\cdot p\cdot f_{1}(p)\,{\mathrm{d}}p-\int_{0}^{1}\gamma(p)\cdot\left(\int_{0}^{1}q\cdot\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p-\int_{0}^{1}\int_{0}^{1}\beta(p,q)\cdot\pi(p,q)\,{\mathrm{d}}q{\mathrm{d}}p$
$\quad+\int_{0}^{1}\alpha_{1}(p)\cdot\left(f_{1}(p)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)\cdot\left(f_{2}(q)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}p\right)\,{\mathrm{d}}q$
$=\int_{0}^{1}\int_{0}^{1}\pi(p,q)\cdot\left(-\beta(p,q)-\gamma(p)\cdot q-\alpha_{1}(p)-\alpha_{2}(q)\right)\,{\mathrm{d}}q{\mathrm{d}}p$
$\quad+\int_{0}^{1}\left(\gamma(p)\cdot p+\alpha_{1}(p)\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)\cdot f_{2}(q)\,{\mathrm{d}}q$
By the first-order condition, the optimal Lagrange multipliers $\beta^{*},\alpha_{1}^{*},\alpha_{2}^{*},\gamma^{*}$ satisfy
$-\beta^{*}(p,q)-\gamma^{*}(p)\cdot q-\alpha_{1}^{*}(p)-\alpha_{2}^{*}(q)=0$
for all $p,q\in[0,1]$. In addition, under the optimal coupling $\pi^{*}$, for any $p,q\in[0,1]$, we have the following complementary slackness conditions:
$\alpha_{1}^{*}(p)+\alpha_{2}^{*}(q)$ $=-\gamma^{*}(p)\cdot q,\quad\pi^{*}(p,q)>0$
$\alpha_{1}^{*}(p)+\alpha_{2}^{*}(q)$ $\leq-\gamma^{*}(p)\cdot q,\quad\pi^{*}(p,q)=0~{}.$
With the above first-order conditions and complementary slackness, the Lagrange dual then becomes
$\mathcal{L}=\int_{0}^{1}\left(\gamma^{*}(p)\cdot p+\alpha_{1}^{*}(p)\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}^{*}(q)\cdot f_{2}(q)\,{\mathrm{d}}q~{},$
which is essentially the objective function in program 11 and its constraints are the complementary slackness conditions that we derive above. This completes the proof of Lemma 3.4. ∎
Proof of Theorem 3.1.
Combining Lemmas 3.2, 3.3 and 3.4 finishes the proof of Theorem 3.1. ∎
3.2 Structure Characterizations of $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$
In this section, we provide additional implications and intuitions behind our proposed relaxed earth mover’s distance $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$.
SCDF representation.  Our proposed relaxed earth mover’s distance $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ admits a succinct closed-form expression in terms of the super-cumulative distribution functions (SCDF).
Proposition 3.5 (SCDF representation).
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$ and the same mean $\mathbb{E}_{p\sim f_{1}}\!\left[{p}\right]=\mathbb{E}_{p\sim f_{2}}\!\left[{p}\right]$, the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ satisfies
$\textsc{REMD}\!\left[{f_{1},f_{2}}\right]=2\cdot\max\nolimits_{t\in[0,1]}~{}S_{1}(t)-S_{2}(t)$
where $S_{1}(t)\triangleq\int_{0}^{t}F_{1}(p)\,{\mathrm{d}}p$ and $S_{2}(t)\triangleq\int_{0}^{t}F_{2}(p)\,{\mathrm{d}}p$ are the SCDFs of distributions $f_{1},f_{2}$, respectively.
Combining the proposition above with Theorem 3.1, we immediately obtain the following corollary for the informativeness gap between perfectly predictors. See Figure 1 for an illustration.
Corollary 3.6.
Given any two perfectly calibrated predictors $\mu,\nu$, informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=2\cdot\max\nolimits_{t\in[0,1]}~{}S_{\mu}(t)-S_{\nu}(t)$
where $S_{\mu}$ and $S_{\nu}$ are the SCDF of prediction distributions $f_{\mu}$ and $f_{\nu}$, respectively.
Proof.
For perfectly calibrated predictors $\mu$ and $\nu$, their prediction distributions has the same mean and support contained by $[0,1]$. Hence, invoking Theorem 3.1 and Proposition 3.5 finishes the proof of Corollary 3.6. ∎
$\frac{1}{2}\cdot\cdot\textsc{InfoGap}\!\left[{\mu,\nu}\right]$$t^{*}$
Figure 1: The gray solid line is $S_{\mu}$ and blue solid line is $S_{\nu}$ where the predictors $\mu$ and $\nu$ are not dominated (in the Blackwell’s sense) with each other. The black dashed line is $S_{h}$ where $h=\mu\vee\nu$ is the join of the predictors $\mu$ and $\nu$, and $t^{*}=\operatorname{argmax}\nolimits_{t\in[0,1]}~{}S_{\mu}(t)-S_{\nu}(t)$. The informativeness gap $\frac{1}{2}\cdot\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ is exactly the height of vertical black solid line.
We remark that Corollary 3.6 together with Proposition 2.4 recovers the SCDF representation of the Blackwell informativeness (Lemma 2.2). We now prove Proposition 3.5.
Proof of Proposition 3.5.
[…]
Define auxiliary function $\omega_{f_{1},f_{2}}:[0,1]\rightarrow\mathbb{R}$ as
$\forall t\in[0,1]:\qquad\omega_{f_{1},f_{2}}(t)\triangleq\displaystyle\int_{0}^{t}f_{1}(p)-f_{2}(p)\,{\mathrm{d}}p$
The objective function of program $\textsc{P}_{1}[f_{1},f_{2}]$ can be reformulated as follows:
$\int_{0}^{1}U(p)\cdot\left(f_{1}(p)-f_{2}(p)\right)~{}{\mathrm{d}}p$ $=\left[U(p)\cdot\omega_{f_{1},f_{2}}(p)\right]_{0}^{1}-\int_{0}^{1}\omega_{f_{1},f_{2}}(p)\,{\mathrm{d}}U(p)=-\int_{0}^{1}\omega_{f_{1},f_{2}}(p)\cdot\delta_{U}(p)\,{\mathrm{d}}p$
where the first equality holds due to integration by parts, and the second equality considers derivative $\delta_{U}(p)\triangleq U^{\prime}(p)$ which exists almost everywhere and uses the fact that $\omega_{f_{1},f_{2}}(0)=\omega_{f_{1},f_{2}}(1)=0$. Consequently, program $\textsc{P}_{1}[f_{1},f_{2}]$ is equivalent to the following program:
$\begin{array}[]{rlll}\inf\limits_{\delta_{U}:[0,1]\rightarrow\mathbb{R}}{}&\displaystyle\int_{0}^{1}\delta_{U}(p)\cdot\omega_{f_{1},f_{2}}(p)~{}{\mathrm{d}}p&\text{s.t.}&\vspace{1mm}\\ &\displaystyle\delta_{U}(p)\in[-1,1],&p\in[0,1]\vspace{1mm}\\ &\displaystyle\delta_{U}(p)\leq\delta_{U}(q)&p\in[0,1],q\in[p,1]\vspace{1mm}\end{array}$ (16)
where the constraints is the sufficient and necessary condition for the derivative $\delta_{U}$ of 1-Lipschitz convex interim utility function $\delta_{U}$.
According to Skreta (2006); Börgers (2015), the extreme points of the uniformly bounded, non-decreasing functions are step functions with only one jump. Thus, the optimal solution $\delta_{U}^{*}$ to the (infinite-dimensional) linear program 16 must be a step function where
$\delta_{U}^{*}(p)=-1\cdot{\mathbbm{1}\left\{p\leq t^{*}\right\}}+1\cdot{\mathbbm{1}\left\{p>t^{*}\right\}}$
for some threshold $t^{*}\in[0,1]$.
With the above observation, to solve program $\textsc{P}_{1}[f_{1},f_{2}]$, it suffices to optimize over all V-shaped functions across different kinks $t\in[0,1]$. In particular, we can write the objective in program $\textsc{P}_{1}[f_{1},f_{2}]$ under a V-shaped function with kink $t\in[0,1]$ as follows:
$\int_{0}^{t}-(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p+\int_{t}^{1}(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p$
[…]
Here the second equality holds since $\int_{0}^{1}(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p=0$, which is implied by the statement assumption that distributions $f_{1},f_{2}$ are valid distribution and have the same mean.
Observe that
$\frac{\partial}{\partial t}\int_{0}^{t}-2(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p$ $=\int_{0}^{t}2\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p-2(t-t)(f_{1}(t)-f_{2}(t))$
$=2(F_{1}(t)-F_{2}(t))$
Thus, the optimal objective value of program $\textsc{P}_{1}[f_{1},f_{2}]$ can be further reformulated as
$\max\nolimits_{t\in[0,1]}~{}\int_{0}^{t}2(F_{1}(p)-F_{2}(p))\,{\mathrm{d}}p=2\cdot\max\nolimits_{t\in[0,1]}~{}S_{1}(t)-S_{2}(t)$
where the equality holds due to SCDF definition. This completes the proof of Proposition 3.5. ∎
Lattice representation.  Given any two distributions $f_{1},f_{2}$ with the same mean and supports contained in $[0,1]$, define their join $f_{\text{join}}\triangleq f_{1}\vee f_{2}$ as the unique distribution whose SCDF $S_{\text{join}}$ satisfies
$\forall p\in[0,1]:\qquad S_{\text{join}}(p)=\max\{S_{1}(p),S_{2}(p)\}$
As a sanity check, since both distributions’ SCDF $S_{1},S_{2}$ are convex, SCDF $S_{\text{join}}$ is also convex and thus $f_{\text{join}}$ is a valid distribution and has the same mean as distributions $f_{1},f_{2}$. As Lemma 2.3 states, the space of information structures (perfectly calibrated predictors) and the Blackwell order over it forms a lattice. Thus, by viewing distributions $f_{1},f_{2}$ and $f_{\text{join}}$ as information structures, join $f_{\text{join}}$ is the least Blackwell informative information structure (perfectly calibrated predictor) that Blackwell-dominants both information structures $f_{1}$ and $f_{2}$.
We now establish the property of our proposed relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ and their corresponding join $f_{\text{join}}=f_{1}\vee f_{2}$.
Proposition 3.7 (Lattice representation).
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$ and the same mean $\mathbb{E}_{p\sim f_{1}}\!\left[{p}\right]=\mathbb{E}_{p\sim f_{2}}\!\left[{p}\right]$, the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ satisfies
$\textsc{REMD}\!\left[{f_{1},f_{2}}\right]=\textsc{REMD}\!\left[{f_{\text{join}},f_{2}}\right]$
where $f_{\text{join}}=f_{1}\vee f_{2}$ is the join of distributions $f_{1},f_{2}$ under the Blackwell order.
Proof.
We rewrite $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ as
$\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ $=2\cdot\max\nolimits_{t\in[0,1]}~{}S_{1}(t)-S_{2}(t)=2\cdot\max\nolimits_{t\in[0,1]}~{}\max\{S_{1}(t),S_{2}(t)\}-S_{2}(t)$
$=2\cdot\max\nolimits_{t\in[0,1]}~{}S_{\text{join}}(t)-S_{2}(t)=\textsc{REMD}\!\left[{f_{\text{join}},f_{2}}\right]$
where the first and last equality holds due to Proposition 3.5, the second equality holds since $S_{1}(0)=S_{2}(0)$ and algebra, and the third equality holds due to the definition of join $f_{\text{join}}$. This completes the proof of Proposition 3.7. ∎
Combining the proposition above with Theorem 3.1, we immediately obtain the following corollary for the informativeness gap between perfectly predictors.
Corollary 3.8.
Given any two perfectly calibrated predictors $\mu,\nu$, informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{InfoGap}\!\left[{h,\nu}\right]$
where $h$ is the perfectly calibrated predictor with prediction distribution $f_{h}=f_{\mu}\vee f_{\nu}$.
With a slight abuse of notation, we also denote by $\mu\vee\nu$ the predictor $h$ in Corollary 3.8, referring to it as the join of predictors $\mu$ and $\nu$.
We make the following interpretation from Corollary 3.8: first, suppose that the perfectly calibrated predictor $\nu$ is more Blackwell informative than the perfectly calibrated predictor $\mu$. By definition, their join is exactly the predictor $\nu$, i.e., $\mu\vee\nu=\nu$. Hence, Corollary 3.8 immediately implies that $\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{InfoGap}\!\left[{\nu,\nu}\right]=0$, which recovers Proposition 2.4.
More importantly, Corollary 3.8 also provides insight into cases where perfectly calibrated predictors $\mu$ and $\nu$ are not comparable under the Blackwell order. In particular, it suggests that the maximum payoff advantage of perfectly calibrated predictor $\mu$ over perfectly calibrated predictor $\nu$ is always achieved in a decision task where predictor $\mu$ performs as well as their join predictor (which is always more Blackwell informative than both). See Figure 2 for an illustration.
$h$$\mu$$\nu$$\textsc{InfoGap}\!\left[{h,\mu}\right]$$\textsc{InfoGap}\!\left[{h,\nu}\right]$$\textsc{InfoGap}\!\left[{\mu,\nu}\right]$$\textsc{InfoGap}\!\left[{\nu,\mu}\right]$
Figure 2: A graphic illustration of join $h\triangleq\mu\vee\nu$ over two calibrated predictors $\mu,\nu$. The solid arrow implies a Blackwell’s order between two predictors, while a dashed arrow not necessarily implies a Blackwell’s order. According to Corollary 3.8, we have that $\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{InfoGap}\!\left[{h,\nu}\right]$ and $\textsc{InfoGap}\!\left[{\nu,\mu}\right]=\textsc{InfoGap}\!\left[{h,\mu}\right]$.
4 The Informativeness of Miscalibrated Predictors
In this section, we characterize the informativeness gap between any two (possibly miscalibrated) predictors. The core of our characterization is a generalized notion of $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ to incorporate miscalibrated predictions, which we define it as follows.
Definition 4.1.
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$ and two functions $\kappa_{1},\kappa_{2}:[0,1]\rightarrow[0,1]$, the informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{1},\kappa_{2}}\right]$ is defined as the optimal objective value of the following optimization program:
$\inf\limits_{\pi\in\overline{\Pi}(f_{1},f_{2})}~{}$ $\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q+(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right|\,{\mathrm{d}}p$ (17)
where
$\overline{\Pi}(f_{1},f_{2})\triangleq\left\{\pi\in\Delta([0,1]\times[0,1]):f_{1}(p)-f_{2}(p)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}q+\int_{0}^{1}\pi(q,p)\,{\mathrm{d}}q=0,~{}~{}p\in[0,1]\right\}$
is referred to as the flow set.
With the above definition, the main result in this section is the following dual characterization of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ when predictors $\mu,\nu$ are possibly miscalibrated.
Theorem 4.1.
For any two (possibly miscalibrated) predictors $\mu,\nu$, their informativeness gap satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$
where $f_{\mu},f_{\nu}$ (resp. $\kappa_{\mu},\kappa_{\nu}$) are the prediction distributions (resp. true expected outcome functions) of predictors $\mu,\nu$, respectively.
The formal proof of Theorem 4.1 is presented in Section 4.1. We discuss the connection between our proposed informativeness gap and other calibration measure in Section 4.2. Below, we make the several remarks regarding Theorem 4.1.
First, unlike our proposed relaxed earth mover’s distance $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$, which is defined over distributions, the new measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ is defined over distributions and two additional functions. This distinction is intuitive, as a miscalibrated predictor $\mu$ is jointly determined by its prediction distribution $f_{\mu}$ and its true expected outcome function $\kappa_{\mu}$. Therefore, a measure that only takes prediction distributions as input does not collect enough information to capture the decision payoff for general miscalibrated predictors.
Second, the flow set $\overline{\Pi}(\cdot,\cdot)$ in the definition above can be viewed as the flow conservation constraint over the 2-dimensional space $[0,1]\times[0,1]$. It is conceptually similar to the flow constraint in the mechanism design literature (Che et al., 2013; Cai et al., 2016). In particular, both our flow constraint and the flow constraint in mechanism design can be interpreted as the dual constraint for interim utility under the best response (or equilibrium). [^9] We conjecture that there might be a deeper connection between our model and the mechanism design model, which we leave for future work.
Third, when the predictors $\mu$ and $\nu$ are both perfectly calibrated, the objective function in program (17) is the same as that in program (4), while the flow set $\overline{\Pi}(f_{\mu},f_{\nu})$ is a strict superset of the coupling set $\Pi(f_{\mu},f_{\nu})$. Theorem 4.1 implies that both $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ and $\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]$ achieve the identical optimal objective value when predictors $\mu$ and $\nu$ are perfectly calibrated. Alternatively, we can summarize this observation as the following characterization for $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$:
Corollary 4.2 (Flow representation).
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$ and the same mean $\mathbb{E}_{p\sim f_{1}}\!\left[{p}\right]=\mathbb{E}_{p\sim f_{2}}\!\left[{p}\right]$, the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]$ satisfies
$\textsc{REMD}\!\left[{f_{1},f_{2}}\right]=\inf\nolimits_{\pi\in\overline{\Pi}(f_{1},f_{2})}\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p~{}.$
where $\overline{\Pi}(f_{1},f_{2})$ is a strict superset of $\Pi(f_{1},f_{2})$ used in the definition of $\textsc{REMD}\!\left[{\cdot,\cdot}\right]$ (Definition 3.2).
Proof.
Since both distributions have supports contained in $[0,1]$ and have the same mean, they can be considered as the prediction distributions of two perfectly calibrated predictors, denoted by $\mu,\nu$. Invoking Theorems 3.1 and 4.1, we obtain $\textsc{REMD}\!\left[{f_{1},f_{2}}\right]=\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{C},\kappa_{C}}\right]$ where function $\kappa_{C}(p)\triangleq p$ for all $p\in[0,1]$. This completes the proof of Corollary 4.2. ∎
Finally, with Corollary 4.2 in mind, one tempting thought is whether directly optimizing over the coupling set $\Pi(f_{\mu},f_{\nu})$ in $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ would suffice–or at least yield a good approximation—-of the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$. However, this is not the case: there exist predictors $\mu,\nu$ for which the gap between optimizing over $\Pi(f_{\mu},f_{\nu})$ in $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$ and $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ can be arbitrarily large.
Example 4.2.
Consider perfectly calibrated predictor $\mu$ which outputs prediction $p=0$ and $p=1$ both with half probability, and miscalibrated predictor $\nu$ with prediction distribution
$f_{\nu}(p)\triangleq\begin{cases}0,&p\in[0,\varepsilon]\\ \displaystyle\frac{\varepsilon}{2(1-2\varepsilon)}\cdot\frac{1}{p^{2}}&p\in(\varepsilon,0.5]\\ \displaystyle\frac{\varepsilon}{2(1-2\varepsilon)}\cdot\frac{1}{(1-p)^{2}}&p\in[0.5,1-\varepsilon)\\ 0,&p\in[1-\varepsilon,1]\\ \end{cases}$
and true expected outcome function $\kappa_{\nu}(p)={\mathbbm{1}\left\{p\geq 0.5\right\}}$.
Proposition 4.3.
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$ and two functions $\kappa_{1},\kappa_{2}:[0,1]\rightarrow[0,1]$, define a generalized version of relaxed EMD, denoted by $\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{1},f_{2},\kappa_{1},\kappa_{2}\right]$ as the optimal objective value of the following optimization program:
$\inf\limits_{\pi\in\Pi(f_{1},f_{2})}\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q+(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right|\,{\mathrm{d}}p~{}.$
where decision variables $\pi$ is optimized over coupling set $\Pi(f_{1},f_{2})$.
For every $\varepsilon\in(0,0.5]$, in the setting of Example 4.2 where predictor $\mu$ is perfectly calibrated while predictor $\nu$ is miscalibrated, it satisfies
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]=\Theta(\varepsilon)\;\;\mbox{and}\;\;\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]=\Omega\left(\varepsilon\ln\frac{1}{\varepsilon}\right)$
and thus
$\frac{\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]}{\textsc{InfoGap}\!\left[{\mu,\nu}\right]}=\Omega\left(\ln\frac{1}{\varepsilon}\right)~{}.$
We remark that the prediction distribution $f_{\nu}$ of the miscalibrated predictor $\nu$, constructed in Example 4.2, is essentially the (truncated) equal-revenue distribution, which is widely used in the mechanism design literature. In addition, for this particular example, the closed-form expressions of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ and $\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]$ are essentially the revenue and welfare in the context of mechanism design, whose ratio is known to be unbounded (Hartline, 2016). Before proving Proposition 4.3, we first present the following lemma, which generalizes Proposition 3.5 (essentially from perfectly calibrated predictors to miscalibrated predictors). Its proof is similar to the one for Proposition 3.5 and can be found in Section A.3.
Lemma 4.4 (Generalized SCDF representation).
Given any two distributions $f_{1},f_{2}$ with supports $\mathsf{supp}(f_{1})\subseteq[0,1],\mathsf{supp}(f_{2})\subseteq[0,1]$ and two functions $\kappa_{1},\kappa_{2}:[0,1]\rightarrow[0,1]$, if $\mathbb{E}_{p\sim f_{1}}\!\left[{\kappa_{1}(p)}\right]=\mathbb{E}_{p\sim f_{2}}\!\left[{\kappa_{2}(p)}\right]$, the informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{1},\kappa_{2}}\right]$ satisfies
$\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{1},\kappa_{2}}\right]=2\cdot\max_{t\in[0,1]}~{}\left(S_{1}(t)+\int_{0}^{t}(p-\kappa_{1}(p))\cdot f_{1}(p)\,{\mathrm{d}}p\right)-\left(S_{2}(t)+\int_{0}^{t}(p-\kappa_{2}(p))\cdot f_{2}(p)\,{\mathrm{d}}p\right)$
where $S_{1}(t)\triangleq\int_{0}^{t}F_{1}(p)\,{\mathrm{d}}p$ and $S_{2}(t)\triangleq\int_{0}^{t}F_{2}(p)\,{\mathrm{d}}p$ are the SCDFs of distributions $f_{1},f_{2}$, respectively.
We now prove Proposition 4.3.
Proof of Proposition 4.3.
Our first step is to lower bound the value $\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]$. In particular, we show that by analyzing the dual of the optimization program in $\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]$, we can lower bound its value as follows  [^10] Notably, two terms $4\int_{0}^{\frac{1}{2}}t\cdot f_{\nu}(t)\,{\mathrm{d}}t$ and $t\cdot\left(1-\frac{t-\varepsilon}{(1-2\varepsilon)t}\right)=t\cdot(1-2\int_{0}^{t}f_{\mu}(p)\,{\mathrm{d}}p)$ can be interpreted as the welfare and revenue in the context of mechanism design.
$\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]\geq 4\int_{0}^{\frac{1}{2}}t\cdot f_{\nu}(t)\,{\mathrm{d}}t=4\int_{0}^{\frac{1}{2}}t\cdot\frac{\varepsilon}{1-2\varepsilon}\cdot\frac{1}{t^{2}}\,{\mathrm{d}}t=\frac{2\varepsilon}{1-2\varepsilon}\ln\frac{1}{2\varepsilon}~{}.$
The above lower bound analysis utilizes the previous analysis in Theorem 3.1 for optimizing over the coupling set $\Pi(f_{\mu},f_{\nu})$ and the specific structure of constructed predictions $\mu,\nu$. We summarize this lower bound in Lemma A.1 and defer it, together with its proof, to Section A.4. Invoking Theorem 4.1 and Lemma 4.4, informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ can be computed as10
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ $=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$
$=2\cdot\max_{t\in[0,1]}~{}\left(S_{\mu}(t)+\int_{0}^{t}(p-\kappa_{\mu}(p))\cdot f_{\mu}(p)\,{\mathrm{d}}p\right)-\left(S_{\nu}(t)+\int_{0}^{t}(p-\kappa_{\nu}(p))\cdot f_{\nu}(p)\,{\mathrm{d}}p\right)$
$=2\cdot\max_{t\in[0,0.5]}~{}\left(S_{\mu}(t)+\int_{0}^{t}(p-\kappa_{\mu}(p))\cdot f_{\mu}(p)\,{\mathrm{d}}p\right)-\left(S_{\nu}(t)+\int_{0}^{t}(p-\kappa_{\nu}(p))\cdot f_{\nu}(p)\,{\mathrm{d}}p\right)$
$=\max_{t\in[0,0.5]}~{}t-\int_{\varepsilon}^{t}\int_{\varepsilon}^{s}\frac{\varepsilon}{1-2\varepsilon}\cdot\frac{1}{p^{2}}\,{\mathrm{d}}p\,{\mathrm{d}}s-\int_{\varepsilon}^{t}p\cdot\frac{\varepsilon}{1-2\varepsilon}\cdot\frac{1}{p^{2}}\,{\mathrm{d}}p$
$=\max_{t\in[0,0.5]}~{}t\cdot\left(1-\frac{t-\varepsilon}{(1-2\varepsilon)t}\right)=\Theta(\varepsilon)$
where the third equality holds since both predictors are symmetric around $t=0.5$.
Combing the two pieces together, we completes the proof of Proposition 4.3. ∎
4.1 Proof of Theorem 4.1
Similar to the proof of Theorem 3.1, we first reformulate the program (3) as a program with treating the univariate form of indirect utility as the variables.
Lemma 4.5.
For any two perfectly calibrated predictors $\mu,\nu$, their informativeness gap satisfies
[…]
$\begin{array}[]{rll}\sup\limits_{U,\delta:[0,1]\rightarrow\mathbb{R}}{}&\displaystyle\int_{0}^{1}U(p)\cdot\left(f_{1}(p)-f_{2}(p)\right)+\delta(p)\cdot\left((\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)\,{\mathrm{d}}p\vspace{1mm}&\text{s.t.}\\ &\delta(p)\in[-1,1]&p\in[0,1]\vspace{1mm}\\ &\displaystyle U(p)-U(q)\leq\delta(p)\cdot(p-q)&p,q\in[0,1]\vspace{1mm}\end{array}$ (21)
Proof.
Invoking Lemma 2.1 and the definition of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$, the objective function in program (3) can be rewritten as follows:
$\int_{0}^{1}U(\kappa_{\mu}(p);p)\cdot f_{\mu}(p)-U(\kappa_{\nu}(p);p)\cdot f_{\nu}(p)~{}{\mathrm{d}}p$
[…]
where function $\delta(p)\triangleq\tilde{u}(p,1)-\tilde{u}(p,0)$. Invoking Lemma 2.5 and the definition of $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$, both constraints are satisfied due to the 1-Lipschitzness and convexity of interim utility $U$, respectively. This complete the proof of Lemma 4.5. ∎
Proof of Theorem 4.1.
We consider the following Lagrange of program $\textsc{P}^{\textsc{MisC}}_{1}[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}]$ with Lagrange multipliers $\lambda:[0,1]\times[0,1]\rightarrow\mathbb{R}_{+}$,
$\mathcal{L}$ $=\int_{0}^{1}U(p)\cdot(f_{\mu}(p)-f_{\nu}(p))~{}{\mathrm{d}}p+\int_{0}^{1}(\kappa_{\mu}(p)-p)\cdot\delta(p)\cdot f_{\mu}(p)~{}{\mathrm{d}}p-\int_{0}^{1}(\kappa_{\nu}(p)-p)\cdot\delta(p)\cdot f_{\nu}(p)~{}{\mathrm{d}}p$
$\quad+\int_{0}^{1}\int_{0}^{1}\lambda(p,q)\cdot\left(\delta(p)\cdot(p-q)-(U(p)-U(q))\right)~{}{\mathrm{d}}p{\mathrm{d}}q$
$=\int_{0}^{1}U(p)\cdot\left(f_{\mu}(p)-f_{\nu}(p)-\int_{0}^{1}\lambda(p,q)\,{\mathrm{d}}q+\int_{0}^{1}\lambda(q,p)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p$
$\quad+\int_{0}^{1}(\kappa_{\mu}(p)-p)\cdot\delta(p)\cdot f_{\mu}(p)~{}{\mathrm{d}}p-\int_{0}^{1}(\kappa_{\nu}(p)-p)\cdot\delta(p)\cdot f_{\nu}(p)~{}{\mathrm{d}}p+\int_{0}^{1}\int_{0}^{1}\lambda(p,q)\cdot\delta(q)\cdot(q-p)~{}{\mathrm{d}}p{\mathrm{d}}q$
By the first-order condition, the optimal dual variable $\lambda^{*}$ satisfies
$f_{\mu}(p)-f_{\nu}(p)-\int_{0}^{1}\lambda^{*}(p,q)\,{\mathrm{d}}q+\int_{0}^{1}\lambda^{*}(q,p)\,{\mathrm{d}}q=0$ (22)
for all $p\in[0,1]$. Thus, under optimal dual variable $\lambda^{*}$, we can simplify $\mathcal{L}$ as follows:
$\mathcal{L}$ $=\int_{0}^{1}\delta(p)\cdot\left(\int_{0}^{1}\lambda^{*}(p,q)\cdot(p-q)~{}{\mathrm{d}}q\right){\mathrm{d}}p+\int_{0}^{1}(\kappa_{\mu}(p)-p)\cdot\delta(p)\cdot f_{\mu}(p)~{}{\mathrm{d}}p-\int_{0}^{1}(\kappa_{\nu}(p)-p)\cdot\delta(p)\cdot f_{\nu}(p)~{}{\mathrm{d}}p$
$=\int_{0}^{1}\delta(p)\cdot\left(\int_{0}^{1}\lambda^{*}(p,q)\cdot(p-q)~{}{\mathrm{d}}q+(\kappa_{\mu}(p)-p)\cdot f_{\mu}(p)-(\kappa_{\nu}(p)-p)\cdot f_{\nu}(p)\right){\mathrm{d}}p$
Since $\delta(p)\in[-1,1]$, under optimal $\delta^{*}$, it satisfies
$\mathcal{L}=\int_{0}^{1}\left|\int_{0}^{1}\lambda^{*}(p,q)\cdot(p-q)~{}{\mathrm{d}}q+(\kappa_{\mu}(p)-p)\cdot f_{\mu}(p)-(\kappa_{\nu}(p)-p)\cdot f_{\nu}(p)\right|{\mathrm{d}}p~{},$
which is equivalent to the objective function in program (17). The flow constraint in program (17) is equivalent to the first-order condition that we derived in Equation (22). Thus, due to Lemma 4.5, program (17) is indeed the dual of program $\textsc{P}^{\textsc{MisC}}_{1}[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}]$. This completes the proof of Theorem 4.1. ∎
4.2 On the Relation to Other Calibration Measures
Our proposed informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ has connections to the existing calibration measures, such as the Expected Calibration Error (ECE) (Foster and Vohra, 1998), and U-Calibration (UCal) (Kleinberg et al., 2023), calibration decision loss (CDL) (Hu and Wu, 2024), distance to calibration (Błasiok et al., 2023). We discuss and summarize these connections below.
Subsuming existing decision-theoretic calibration measures.  We first show that the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ subsumes existing decision-theoretic calibration measures.
Definition 4.3 (U-Calibration (UCal), adopted from Kleinberg et al., 2023).
Given any (possibly miscalibrated) predictor $\mu$, its U-Calibration, denoted by $\textsc{UCal}\!\left[{\mu}\right]$, is defined as follows:
$\textsc{UCal}\!\left[{\mu}\right]\triangleq\sup\nolimits_{(\mathcal{A},u)\in\mathcal{U}}~{}\displaystyle\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu^{\textsc{NI}}}\right]-\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu}\right]~{},$
where $\mu^{\textsc{NI}}$ denotes the perfectly calibrated predictor that consistently generates prediction $\mathbb{E}_{p\sim f_{\mu}}\!\left[{\kappa_{\mu}(p)}\right]$, i.e., its prediction distribution $f_{\mu^{\textsc{NI}}}\triangleq\textsc{Dirac}(\mathbb{E}_{p\sim f_{\mu}}\!\left[{\kappa_{\mu}(p)}\right])$ where Dirac is the Dirac delta function.
Definition 4.4 (Calibration decision loss (CDL), adopted from Hu and Wu, 2024).
Given any (possibly miscalibrated) predictor $\mu$, its calibration decision loss, denoted by $\textsc{CDL}\!\left[{\mu}\right]$, is defined as follows:
$\textsc{CDL}\!\left[{\mu}\right]\triangleq\sup\nolimits_{(\mathcal{A},u)\in\mathcal{U}}~{}\displaystyle\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu^{\textsc{True}}}\right]-\textsc{Payoff}_{(\mathcal{A},u)}\!\left[{\mu}\right]~{},$
where $\mu^{\textsc{True}}$ denotes the perfectly calibrated predictor that generates the true expected outcome conditional on the prediction that is generated by the predictor $\mu$.
First, given any miscalibrated predictor $\nu$, we observe that by letting $\mu$ to be some perfectly calibrated predictors that are related to $\nu$, informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ subsumes the UCal and CDL as follows.
Fact 4.6.
Given any miscalibrated predictor $\mu$, we have
$\textsc{InfoGap}\!\left[{\mu^{\textsc{NI}},\mu}\right]=\textsc{UCal}\!\left[{\mu}\right];\quad\textsc{InfoGap}\!\left[{\mu^{\textsc{True}},\mu}\right]=\textsc{CDL}\!\left[{\mu}\right]~{}.$
Upper bounding the informativeness gap.  We next show that we can use existing calibration measures (e.g., the Expected calibration error (ECE), defined as in Definition 4.5), together with our relaxed earth mover’s distance (REMD), to upper bound informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$.
Definition 4.5 (Expected calibration error (ECE)).
Given any (possibly miscalibrated) predictor $\mu$, its expected calibration error, denoted by $\textsc{ECE}\!\left[{\mu}\right]$, is defined as follows:
$\textsc{ECE}\!\left[{\mu}\right]\triangleq\mathbb{E}_{p\sim f_{\mu}}\!\left[{\left|p-\kappa_{\mu}(p)\right|}\right]=\int_{0}^{1}f_{\mu}(p)\cdot\left|p-\kappa_{\mu}(p)\right|\,{\mathrm{d}}p~{}.$
With the above definition, we can obtain the following upper bound for $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$.
Proposition 4.7.
For any two (possibly miscalibrated) predictors $\mu,\nu$, their informativeness gap can be upper bounded as
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]\leq\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]+\textsc{ECE}\!\left[{\mu}\right]+\textsc{ECE}\!\left[{\nu}\right]~{}.$
Proof.
Given any two predictors $\mu,\nu$, let
$\pi_{\textsc{REMD}}^{*}$ $=\arg\inf\nolimits_{\pi\in\Pi(f_{\mu},f_{\nu})}\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p~{}.$
be the optimal coupling that achieves the relaxed earth’s mover distance between the two prediction distributions $f_{\mu}$ and $f_{\nu}$. Clearly we also have $\pi_{\textsc{REMD}}^{*}\in\overline{\Pi}(f_{\mu},f_{\nu})$. From Theorem 4.1, we have
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ $=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$
$=\inf_{\pi\in\overline{\Pi}(f_{\mu},f_{\nu})}\int_{0}^{1}\left|\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q+(\kappa_{\mu}(p)-p)f_{\mu}(p)-(\kappa_{\nu}(p)-p)f_{\nu}(p)\right|\,{\mathrm{d}}p$
$\leq\int_{0}^{1}\left|\int_{0}^{1}\pi_{\textsc{REMD}}^{*}(p,q)\cdot(p-q)\,{\mathrm{d}}q+(\kappa_{\mu}(p)-p)f_{\mu}(p)-(\kappa_{\nu}(p)-p)f_{\nu}(p)\right|\,{\mathrm{d}}p$
$\leq\int_{0}^{1}\left|\int_{0}^{1}\pi_{\textsc{REMD}}^{*}(p,q)\cdot(p-q)\,{\mathrm{d}}q\right|\,{\mathrm{d}}p+\textsc{ECE}\!\left[{\mu}\right]+\textsc{ECE}\!\left[{\nu}\right]$
$=\textsc{REMD}\!\left[{f_{\mu},f_{\nu}}\right]+\textsc{ECE}\!\left[{\mu}\right]+\textsc{ECE}\!\left[{\nu}\right]~{}.$
where the first inequality holds since $\pi_{\textsc{REMD}}^{*}\in\overline{\Pi}(f_{\mu},f_{\nu})$, and the second inequality holds due to the triangle inequality. This completes the proof of Proposition 4.7. ∎
Combing Proposition 4.7 and 4.6, we obtain the following tighter upper bound on CDL.
Corollary 4.8.
Given any (possibly miscalibrated) predictor $\mu$, let $\mu^{\textsc{True}}$ be its perfectly calibrated predictor (see construction in Definition 4.4), its calibration decision loss can be upper bounded as
$\textsc{CDL}\!\left[{\mu}\right]\leq\textsc{REMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]+\textsc{ECE}\!\left[{\mu}\right]\leq 2\textsc{ECE}\!\left[{\mu}\right]~{}.$
Proof.
Given any miscalibrated predictor $\mu$ and its corresponding perfectly calibrated predictor $\mu^{\textsc{True}}$, we know that
$\textsc{CDL}\!\left[{\mu}\right]$ $=\textsc{InfoGap}\!\left[{\mu^{\textsc{True}},\mu}\right]$ (by 4.6)
$=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu},\kappa_{\mu^{\textsc{True}}},\kappa_{\mu}}\right]$ (by Theorem 4.1)
$\leq\textsc{REMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]+\textsc{ECE}\!\left[{\mu^{\textsc{True}}}\right]+\textsc{ECE}\!\left[{\mu}\right]$ (by Proposition 4.7)
$=\textsc{REMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]+\textsc{ECE}\!\left[{\mu}\right]$ ($\textsc{ECE}\!\left[{\mu^{\textsc{True}}}\right]=0$)
$\leq\textsc{EMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]+\textsc{ECE}\!\left[{\mu}\right]$ (by triangle inequality)
$\leq 2\textsc{ECE}\!\left[{\mu}\right]$
which finishes the proof of Corollary 4.8. ∎
Previous work Hu and Wu (2024) have established that $\textsc{CDL}\!\left[{\mu}\right]\leq 2\textsc{ECE}\!\left[{\mu}\right]$. Here, we obtain a slightly tighter upper bound where the CDL of a predictor $\mu$ can be upper bounded by the summation of the relaxed earth mover’s distance $\textsc{REMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]$ between the two prediction distributions $f_{\mu^{\textsc{True}}},f_{\mu}$ and its $\textsc{ECE}\!\left[{\mu}\right]$. Notice that we have
$\textsc{REMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]\leq\textsc{EMD}\!\left[{f_{\mu^{\textsc{True}}},f_{\mu}}\right]\leq\textsc{ECE}\!\left[{\mu}\right]~{}.$
Trilemma for complete, sound, and continuous informativeness measure.  Błasiok et al. (2023) proposed one desiderata for an ideal calibration measure that should satisfy: the consistency – the measure is robust completeness (correct predictions have low error) and robust soundness (incorrect predictions have high error). Our proposed informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ can be served as a tool for quantifying the informativeness gap between any pair of predictors. By Theorem 4.1, $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ is both complete and sound (in fact, it satisfies these criteria exactly), and thus it is consistent.
However, modern machine learning predictors typically make predictions continuously over a probabilistic space. This motivates an additional desideratum: we seek calibration error measures that are continuous in prediction values, meaning they should not be sensitive to small perturbations in predictions. For instance, the distance to calibration measure (or the earth mover’s distance $\textsc{EMD}\!\left[{\cdot,\cdot}\right]$ between predictors) (Błasiok et al., 2023), the smooth calibration error (Kakade and Foster, 2008) and the smooth ECE (Kuleshov et al., 2018), all satisfy this desirable continuity property.
This raises a natural question: Does our informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ also satisfy such a continuity property? If not, is there an alternative informativeness measure, $\textsc{InfoMeasure}\!\left[{\cdot,\cdot}\right]$ [^11] Here we allow the informativeness measure $\textsc{InfoMeasure}\!\left[{\cdot,\cdot}\right]$ to be an arbitrary function over the pair of two predictors., that is both consistent w.r.t. informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ and “continuous”?
Unfortunately, as we show below, our informativeness measure $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{\cdot,\cdot,\cdot,\cdot}\right]$ fails to satisfy a standard notion of continuity (specifically, continuity under the earth mover’s distance $\textsc{EMD}\!\left[{\cdot,\cdot}\right]$). [^12] We adopt $\textsc{EMD}\!\left[{\cdot,\cdot}\right]$ as our continuity measure. This is also widely adopted in the literature (e.g., Błasiok et al., 2023; Hartline et al., 2025). Moreover, we establish a stronger impossibility result: no informativeness measure $\textsc{InfoMeasure}\!\left[{\cdot,\cdot}\right]$ can satisfy completeness, soundness, and continuity simultaneously.
Proposition 4.9 (Informativeness measure trilemma).
For any informativeness measure $\textsc{InfoMeasure}\!\left[{\cdot,\cdot}\right]$ that takes as input any pair of predictors, at least one of the following property must fail:
  • (Approximate completeness): for any predictors $\mu,\nu$,
    $\textsc{InfoMeasure}\!\left[{\mu,\nu}\right]=O\left(\left(\textsc{InfoGap}\!\left[{\mu,\nu}\right]\right)^{c}\right)$
    where $c>0$ is an absolute constant.
  • (Approximate soundness): for any predictors $\mu,\nu$,
    $\textsc{InfoMeasure}\!\left[{\mu,\nu}\right]=\Omega\left(\left(\textsc{InfoGap}\!\left[{\mu,\nu}\right]\right)^{s}\right)$
    where $s>0$ is an absolute constant.
  • (Continuity): for any predictors $\mu,\mu^{\dagger},\nu,\nu^{\dagger}$,
    $\left|\textsc{InfoMeasure}\!\left[{\mu,\nu}\right]-\textsc{InfoMeasure}\!\left[{\mu^{\dagger},\nu^{\dagger}}\right]\right|=O\left(\textsc{EMD}\!\left[{f_{\mu},f_{\mu^{\dagger}}}\right]+\textsc{EMD}\!\left[{f_{\nu},f_{\nu^{\dagger}}}\right]\right)$
    where $\textsc{EMD}\!\left[{\cdot,\cdot}\right]$ is defined as in Definition 3.1.
Proof.
We prove the proposition statement by contradiction. Suppose there exists informativeness measure $\textsc{InfoMeasure}\!\left[{\cdot,\cdot}\right]$ that satisfies all three properties. Fix a sufficiently small $\varepsilon>0$. We consider the following two predictors:
  • Predictor $\mu$ is perfectly calibrated with a point mass on $p=0.5$;
  • Predictor $\nu$ is miscalibrated and satisfies that $f_{\nu}(0.5+\varepsilon)=f_{\nu}(0.5-\varepsilon)=0.5$, and its true expected outcome function satisfies $\kappa_{\nu}(0.5+\varepsilon)=0$, and $\kappa_{\nu}(0.5-\varepsilon)=1$.
By construction, the informativeness gap satisfies
$\textsc{InfoGap}\!\left[{\mu,\mu}\right]$ $=0$
and
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ $=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$
$=2\cdot\max_{t\in[0,1]}~{}\left(S_{\mu}(t)+\int_{0}^{t}(p-\kappa_{\mu}(p))\cdot f_{\mu}(p)\,{\mathrm{d}}p\right)-\left(S_{\nu}(t)+\int_{0}^{t}(p-\kappa_{\nu}(p))\cdot f_{\nu}(p)\,{\mathrm{d}}p\right)$
$=\Theta(1)$
where the first equality holds due to Theorem 4.1, the second equality holds due to Lemma 4.4, and the third equality holds by algebra. Due to the approximate completeness property and the approximate soundness property, the informativeness measure $\textsc{InfoMeasure}\!\left[{\cdot,\cdot}\right]$ satisfies
$\textsc{InfoMeasure}\!\left[{\mu,\mu}\right]=0\;\;\mbox{and}\;\;\textsc{InfoMeasure}\!\left[{\mu,\nu}\right]=\Theta(1)$
Therefore,
$\Theta(1)=\left|\textsc{InfoMeasure}\!\left[{\mu,\nu}\right]-\textsc{InfoMeasure}\!\left[{\mu,\mu}\right]\right|=O(\textsc{EMD}\!\left[{f_{\mu},f_{\mu}}\right]+\textsc{EMD}\!\left[{f_{\nu},f_{\mu}}\right])=O(\varepsilon)$
which is a contradiction. (Here the second equality holds due to the continuity property.) This completes the proof of Proposition 4.9. ∎
5 Sample Efficiency for Estimating $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$
In practice, we may not have direct access to the prediction distribution $f_{\mu}$ or the true expected outcome function $\kappa_{\mu}$ of a given predictor $\mu$. Instead, we often only have sample access. Motivated by this, we study in this section the sample complexity of estimating informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$. Our main result shows that informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ can be estimated sample-efficiently under the prediction-only access model, defined as follows: [^13] In addition to the prediction-only access model, prior work (Dwork et al., 2021; Błasiok et al., 2023) also considers a more powerful alternative, known as the sample-access model, which allows sampling of feature–prediction–label triples. We also remark that our proposed informativeness gap $\textsc{InfoGap}\!\left[{\cdot,\cdot}\right]$ is inherently defined under the prediction-only access model. This stands in contrast to the ground-truth distance from calibration proposed in Błasiok et al. (2023), which is defined under the more powerful sample-access model.
Definition 5.1 (Prediction-only access model, adopted from Błasiok et al., 2023).
Given a (possibly miscalibrated) predictor $\mu$, the prediction-only access model assumes the existence of an oracle that samples prediction-label pairs $(p,y)$ consistent with predictor $\mu$. Specifically, prediction $p$ is drawn from the prediction distribution $f_{\mu}$, and conditional on $p$, the outcome $y$ is drawn from a Bernoulli distribution $\mathsf{Bern}(\kappa_{\mu}(p))$ with mean equal to the true expected outcome $\kappa_{\mu}(p)$.
Theorem 5.1 (Sample complexity).
Fix any two (possibly miscalibrated) predictors $\mu,\nu$ and any $\varepsilon,\delta\in(0,1)$. Define $n\triangleq\Theta(\frac{1}{\varepsilon^{2}}\ln\frac{1}{\delta})$. With $n$ i.i.d. prediction-label samples $\{(p_{i}^{\mu},y_{i}^{\mu})\}_{i\in[n]}$ from predictor $\mu$ and $n$ i.i.d. prediction-label samples $\{(p_{i}^{\nu},y_{i}^{\nu})\}_{i\in[n]}$ from predictor $\nu$, define estimator $\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]$ as: [^14] Operator ${\left(\cdot\right)^{+}}\triangleq\max\{0,\cdot\}$.
$\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]\triangleq\max\limits_{t\in[0,1]}~{}\frac{1}{n}\sum\limits_{i\in[n]}\left({\left(t-p_{i}^{\mu}\right)^{+}}+(p_{i}^{\mu}-y_{i}^{\mu})\cdot{\mathbbm{1}\left\{\mu_{i}^{\mu}\leq t\right\}}\right)-\left({\left(t-p_{i}^{\nu}\right)^{+}}+(p_{i}^{\nu}-y_{i}^{\nu})\cdot{\mathbbm{1}\left\{\mu_{i}^{\nu}\leq t\right\}}\right)$
Then, informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ and estimator $\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]$ satisfy
$\left|\textsc{InfoGap}\!\left[{\mu,\nu}\right]-\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]\right|\leq\varepsilon$
with probability at least $1-\delta$.
For a miscalibrated predictor, the true expected outcome function can be highly non-smooth and, as a result, is unlikely to be estimated in a sample-efficient manner. However, although this function plays a crucial role in determining the agent’s payoff (e.g., see Lemma 2.1), we bypass this challenge in Theorem 5.1 by avoiding the need to estimate the true expected outcome function directly.
Proof of Theorem 5.1.
Invoking Theorem 4.1 and Lemma 4.4, the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ can be expressed as
$\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ $=\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}}\right]$
$=2\cdot\max_{t\in[0,1]}~{}\left(S_{\mu}(t)+\int_{0}^{t}(p-\kappa_{\mu}(p))\cdot f_{\mu}(p)\,{\mathrm{d}}p\right)-\left(S_{\nu}(t)+\int_{0}^{t}(p-\kappa_{\nu}(p))\cdot f_{\nu}(p)\,{\mathrm{d}}p\right)$
Note that
$S_{\mu}(t)$ $=\int_{0}^{t}F_{\mu}(p)\,{\mathrm{d}}p=\mathbb{E}_{p\sim f_{\mu}}\!\left[{{\left(t-p\right)^{+}}}\right]$
and
$\int_{0}^{t}(p-\kappa_{\mu}(p))\cdot f_{\mu}(p)$ $=\mathbb{E}_{p\sim f_{\mu}}\!\left[{{(p-\kappa_{\mu}(p))\cdot{\mathbbm{1}\left\{p\leq t\right\}}}}\right]$
$=\mathbb{E}_{p\sim f_{\mu}}\!\left[{{\left(p-\mathbb{E}_{y\sim\mathsf{Bern}(\kappa_{\mu}(p))}\!\left[{y}\right]\right)\cdot{\mathbbm{1}\left\{p\leq t\right\}}}}\right]$
$=\mathbb{E}_{p\sim f_{\mu}}\!\left[{\mathbb{E}_{y\sim\mathsf{Bern}(\kappa_{\mu}(p))}\!\left[{(p-y)\cdot{\mathbbm{1}\left\{p\leq t\right\}}}\right]}\right]$
where the last equality holds due to the law of total expectation. Similar expressions hold for predictor $\nu$. Thus, the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$ can be expressed as
$\begin{split}\textsc{InfoGap}\!\left[{\mu,\nu}\right]&=2\cdot\max_{t\in[0,1]}~{}\left(\mathbb{E}_{p\sim f_{\mu}}\!\left[{\mathbb{E}_{y\sim\mathsf{Bern}(\kappa_{\mu}(p))}\!\left[{{\left(t-p\right)^{+}}+(p-y)\cdot{\mathbbm{1}\left\{p\leq t\right\}}}\right]}\right]\right.\\ &\qquad\qquad\qquad\left.-\mathbb{E}_{p\sim f_{\nu}}\!\left[{\mathbb{E}_{y\sim\mathsf{Bern}(\kappa_{\nu}(p))}\!\left[{{\left(t-p\right)^{+}}+(p-y)\cdot{\mathbbm{1}\left\{p\leq t\right\}}}\right]}\right]\right)\end{split}$ (23)
Given the construction of $\widehat{\textsc{InfoGap}}\!\left[\cdot,\cdot\right]$, we conclude that $\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]$ is a consistent estimator of the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$.
We next analyze the sample complexity of estimator $\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]$. Note that in the reformulation (23) of the informativeness gap $\textsc{InfoGap}\!\left[{\mu,\nu}\right]$, the family of functions $\{{\left(t-p\right)^{+}}\}_{t\in[0,1]}$ has a pseudo-dimension of $1$, and family of of functions $\{(p-y)\cdot{\mathbbm{1}\left\{p\leq t\right\}}\}_{t\in[0,1]}$ also has a pseudo-dimension of $1$. Therefore, by the uniform convergence theorem (for bounded real-valued class), we have that when the number of samples, $n$, satisfies
$n\geq\frac{16}{\varepsilon^{2}}\cdot\left(\ln(2en)+\ln\frac{8}{\delta}\right)$
each of the following holds with probability at least $1-\frac{\delta}{2}$:
$\max_{t\in[0,1]}\left|\frac{1}{n}\sum\limits_{i\in[n]}{\left(t-p_{i}^{\mu}\right)^{+}}+(p_{i}^{\mu}-y_{i}^{\mu})\cdot{\mathbbm{1}\left\{\mu_{i}^{\mu}\leq t\right\}}-\mathbb{E}_{p\sim f_{\mu}}\!\left[{\mathbb{E}_{y\sim\mathsf{Bern}(\kappa_{\mu}(p)}\!\left[{{\left(t-p\right)^{+}}+(p-y)\cdot{\mathbbm{1}\left\{p\leq t\right\}}}\right]}\right]\right|$ $\leq\frac{\varepsilon}{2}$
$\max_{t\in[0,1]}\left|\frac{1}{n}\sum\limits_{i\in[n]}{\left(t-p_{i}^{\nu}\right)^{+}}+(p_{i}^{\nu}-y_{i}^{\nu})\cdot{\mathbbm{1}\left\{\mu_{i}^{\nu}\leq t\right\}}-\mathbb{E}_{p\sim f_{\nu}}\!\left[{\mathbb{E}_{y\sim\mathsf{Bern}(\kappa_{\nu}(p)}\!\left[{{\left(t-p\right)^{+}}+(p-y)\cdot{\mathbbm{1}\left\{p\leq t\right\}}}\right]}\right]\right|$ $\leq\frac{\varepsilon}{2}$
Thus, invoking the triangle inequality, we obtain
$\left|\textsc{InfoGap}\!\left[{\mu,\nu}\right]-\widehat{\textsc{InfoGap}}\!\left[\mu,\nu\right]\right|\leq\varepsilon$
and completes the proof of Theorem 5.1. ∎
Appendix A Missing Proofs
In this section, we provide all the missing proofs omitted from the main text.
A.1 Proof of Proposition 2.4
See 2.4
Proof.
The only if direction is trivial since the definition of Blackwell informativeness considers all decision problems.
For the if direction, It suffices to show for all (possibly unbounded) convex (univariate) interim utility functions the expected utility under predictor $\mu$ is weakly higher than the expected utility under predictor $\nu$. To see this, fix an arbitrary (possibly unbounded) convex interim utility function $U$. Define bounded convex interim utility function $U^{(k)}$ which is obtained by truncating the derivative of function $U$ at a large constant $k$ from top. Consider the sequence of bounded convex interim utility function $(U^{(k)})$ as $k$ goes to infinite. Observe that $\textsc{InfoGap}\!\left[{\nu,\mu}\right]=0$ implies that for every bounded convex interim utility function $U^{(k)}$, the expected utility under predictor $\mu$ is weakly higher than the expected utility under predictor $\nu$, i.e., $\mathbb{E}_{p\sim f_{\mu}}\!\left[{U^{(k)}(p)}\right]\geq\mathbb{E}_{p\sim f_{\nu}}\!\left[{U^{(k)}(p)}\right]$. Invoking the dominated convergence theorem, we obtain $\lim_{k\rightarrow\infty}\mathbb{E}_{p\sim f_{\mu}}\!\left[{U^{(k)}(p)}\right]=\mathbb{E}_{p\sim f_{\mu}}\!\left[{U(p)}\right]$ and $\lim_{k\rightarrow\infty}\mathbb{E}_{p\sim f_{\nu}}\!\left[{U^{(k)}(p)}\right]=\mathbb{E}_{p\sim f_{\nu}}\!\left[{U(p)}\right]$. Hence, $\mathbb{E}_{p\sim f_{\mu}}\!\left[{U(p)}\right]\geq\mathbb{E}_{p\sim f_{\nu}}\!\left[{U(p)}\right]$. This completes the proof of Proposition 2.4. ∎
A.2 Proof of Lemma 2.5
See 2.5
Proof.
According to Kleinberg et al. [2023], Hu and Wu [2024], for every decision problem (of the agent), its interim utility function is equivalent to a proper scoring function, which is piecewise linear with maximum absolute slope of $\max_{p\in[0,1]}|\tilde{u}(p,1)-\tilde{u}(p,0)|$. The bounded utility difference $\max_{p\in[0,1]}|\tilde{u}(p,1)-\tilde{u}(p,0)|\leq 1$ is equivalent to the 1-Lipschitzness of the univariate form of the interim utility function $U$. Finally, recall that a scoring function is proper if and only if its univariate form is convex [McCarthy, 1956]. This completes the proof of Lemma 2.5. ∎
A.3 Proof of Lemma 4.4
See 4.4
Proof.
Invoking Theorem 4.1 and Lemma 4.5, we obtain the identity between $\textsc{REMD${}^{\textsc{MisC}}$}\!\left[{f_{1},f_{2},\kappa_{1},\kappa_{2}}\right]$ and the optimal objective of program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$, i.e.,
[…]
Define auxiliary function $\omega_{f_{1},f_{2}}:[0,1]\rightarrow\mathbb{R}$ as
$\forall t\in[0,1]:\qquad\omega_{f_{1},f_{2}}(t)\triangleq\displaystyle\int_{0}^{t}f_{1}(p)-f_{2}(p)\,{\mathrm{d}}p$
The objective function of program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$ can be reformulated as follows:
$\int_{0}^{1}U(p)\cdot\left(f_{1}(p)-f_{2}(p)\right)+\delta(p)\cdot\left((\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)\,{\mathrm{d}}p$
[…]
where the first equality holds due to integration by parts, and the second equality considers derivative $\delta_{U}(p)\triangleq U^{\prime}(p)$ which exists almost everywhere and uses the fact that $\omega_{f_{1},f_{2}}(0)=\omega_{f_{1},f_{2}}(1)=0$. To understand the third equality, note that constraint $U(p)-U(q)\leq\delta(p)\cdot(p-q)$ for all $p,q\in[0,1]$ in program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$ implies that
  • function $\delta$ is non-decreasing. To see this, fix any $0\leq p<q\leq 1$. Observe that for every pair $(p,q)$, $U(p)-U(q)\leq\delta(p)\cdot(p-q)$, while for pair $(q,p)$, $U(q)-U(p)\leq\delta(q)\cdot(q-p)$. Thus, combining these two inequalities we have
    $\delta(p)\leq\frac{U(q)-U(p)}{q-p}\leq\delta(q)~{},$
    which shows the monotonicity of function $\delta$.
  • Furthermore, it guarantees that for every $p\in[0,1]$, $\delta(p)\in[U^{\prime}_{-}(p),U^{\prime}_{+}(p)]$ where $U^{\prime}_{-}(p)$ and $U^{\prime}_{+}(p)$ are the left and right derivative of function $U$ at $p$. This can be verified by plugging in $(p,p-\varepsilon)$ and $(p+\varepsilon,p)$ for sufficiently small $\varepsilon$ in the constraint $U(p)-U(q)\leq\delta(p)\cdot(p-q)$. Since $\delta\in[U^{\prime}_{-}(p),U^{\prime}_{+}(p)]$ and it is monotone, $U$ is a convex function, and thus $U^{\prime}_{-}(p)=U^{\prime}_{+}(q)$ almost everywhere except on a set of Lebesgue measure 00. Thus, functions $\delta_{U}=\delta$ almost everywhere.
Consequently, program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$ is equivalent to the following program:
$\begin{array}[]{rlll}\inf\limits_{\delta_{U}:[0,1]\rightarrow\mathbb{R}}{}&\displaystyle\int_{0}^{1}\delta_{U}(p)\cdot\left(\omega_{f_{1},f_{2}}(p)-(\kappa_{1}(p)-p)\cdot f_{1}(p)+(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)~{}{\mathrm{d}}p&\text{s.t.}&\vspace{1mm}\\ &\displaystyle\delta_{U}(p)\in[-1,1],&p\in[0,1]\vspace{1mm}\\ &\displaystyle\delta_{U}(p)\leq\delta_{U}(q)&p\in[0,1],q\in[p,1]\vspace{1mm}\end{array}$
According to Skreta [2006], Börgers [2015], the extreme points of the uniformly bounded, non-decreasing functions are step functions with only one jump. Thus, the optimal solution $\delta_{U}^{*}$ to the above (infinite-dimensional) linear program must be a step function where
$\delta_{U}^{*}(p)=-1\cdot{\mathbbm{1}\left\{p\leq t^{*}\right\}}+1\cdot{\mathbbm{1}\left\{p>t^{*}\right\}}$
for some threshold $t^{*}\in[0,1]$.
With the above observation, to solve program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$, it suffices to optimize over all V-shaped functions across different kinks $t\in[0,1]$. In particular, we can write the objective in program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$ under a V-shaped function with kink $t\in[0,1]$ as follows:
$\int_{0}^{t}-(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p+\int_{t}^{1}(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p$
$\quad-\int_{0}^{t}(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\,{\mathrm{d}}p+\int_{t}^{1}(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\,{\mathrm{d}}p$
[…]
$\quad-\int_{0}^{t}2\left((\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)\,{\mathrm{d}}p+\int_{0}^{1}(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\,{\mathrm{d}}p$
[…]
Here the second equality holds since $\int_{0}^{1}(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p+\int_{0}^{1}(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\,{\mathrm{d}}p=0$, which is implied by the statement assumption that distributions $f_{1},f_{2}$ are valid distribution and have the same mean.
Observe that
$\frac{\partial}{\partial t}\int_{0}^{t}-2(p-t)\cdot(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p$ $=\int_{0}^{t}2(f_{1}(p)-f_{2}(p))\,{\mathrm{d}}p-2(t-t)\cdot(f_{1}(t)-f_{2}(t))$
$=2(F_{1}(t)-F_{2}(t))$
Thus, the optimal objective value of program $\textsc{P}^{\textsc{MisC}}_{1}[f_{1},f_{2},\kappa_{1},\kappa_{2}]$ can be further reformulated as
$\max\nolimits_{t\in[0,1]}~{}\int_{0}^{t}2(F_{1}(p)-F_{2}(p))\,{\mathrm{d}}p-\int_{0}^{t}2\left((\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)\,{\mathrm{d}}p$
[…]
where the equality holds due to SCDF definition. This completes the proof of Lemma 4.4. ∎
A.4 A Missing Proof in Proposition 4.3
Lemma A.1.
For the predictors $\mu,\nu$ defined as in Example 4.2, we have that $\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]\geq 4\int_{0}^{\frac{1}{2}}t\cdot f_{\nu}(t)\,{\mathrm{d}}t$.
Proof.
We consider Lagrange multipliers $\alpha_{1},\alpha_{2}:[0,1]\rightarrow\mathbb{R}$, and $\beta:[0,1]\times[0,1]\rightarrow\mathbb{R}_{+}$ for the optimization program in Proposition 4.3:
$\mathcal{L}$ $=\int_{0}^{1}\gamma(p)\cdot\left(\int_{0}^{1}\pi(p,q)\cdot(p-q)\,{\mathrm{d}}q+(\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)\,{\mathrm{d}}p$
$\quad-\int_{0}^{1}\int_{0}^{1}\beta(p,q)\cdot\pi(p,q)\,{\mathrm{d}}q{\mathrm{d}}p$
$\quad+\int_{0}^{1}\alpha_{1}(p)\cdot\left(f_{1}(p)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)\cdot\left(f_{2}(q)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}p\right)\,{\mathrm{d}}q$
$=\int_{0}^{1}\gamma(p)\cdot p\cdot f_{1}(p)\,{\mathrm{d}}p-\int_{0}^{1}\gamma(p)\cdot\left(\int_{0}^{1}q\cdot\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p-\int_{0}^{1}\int_{0}^{1}\beta(p,q)\cdot\pi(p,q)\,{\mathrm{d}}q{\mathrm{d}}p$
$\quad+\int_{0}^{1}\gamma(p)\cdot\left((\kappa_{1}(p)-p)\cdot f_{1}(p)-(\kappa_{2}(p)-p)\cdot f_{2}(p)\right)\,{\mathrm{d}}p$
$\quad+\int_{0}^{1}\alpha_{1}(p)\cdot\left(f_{1}(p)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}q\right)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}(q)\cdot\left(f_{2}(q)-\int_{0}^{1}\pi(p,q)\,{\mathrm{d}}p\right)\,{\mathrm{d}}q$
$=\int_{0}^{1}\int_{0}^{1}\pi(p,q)\cdot\left(-\beta(p,q)-\gamma(p)\cdot q-\alpha_{1}(p)-\alpha_{2}(q)\right)\,{\mathrm{d}}q{\mathrm{d}}p$
$\quad+\int_{0}^{1}\left(\gamma(p)\cdot\kappa_{1}(p)+\alpha_{1}(p)\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}(\alpha_{2}(p)-\gamma(p)(\kappa_{2}(p)-p))\cdot f_{2}(p)\,{\mathrm{d}}p$
By the first-order condition, the optimal Lagrange multipliers $\beta^{*},\alpha_{1}^{*},\alpha_{2}^{*},\gamma^{*}$ satisfy
$-\beta^{*}(p,q)-\gamma^{*}(p)\cdot q-\alpha_{1}^{*}(p)-\alpha_{2}^{*}(q)=0$
for all $p,q\in[0,1]$. In addition, under the optimal coupling $\pi^{*}$, for any $p,q\in[0,1]$, we have the following complementary slackness conditions:
$\alpha_{1}^{*}(p)+\alpha_{2}^{*}(q)$ $=-\gamma^{*}(p)\cdot q,\quad\pi^{*}(p,q)>0$
$\alpha_{1}^{*}(p)+\alpha_{2}^{*}(q)$ $\leq-\gamma^{*}(p)\cdot q,\quad\pi^{*}(p,q)=0~{}.$
With the above first-order conditions and complementary slackness, together with the miscalibrated predictions defined as in Example 4.2, the Lagrange dual then becomes
$\mathcal{L}$ $=\int_{0}^{1}\left(\gamma^{*}(p)\cdot\kappa_{1}(p)+\alpha_{1}^{*}(p)\right)\cdot f_{1}(p)\,{\mathrm{d}}p+\int_{0}^{1}(\alpha_{2}^{*}(p)-\gamma^{*}(p)(\kappa_{2}(p)-p))\cdot f_{2}(p)\,{\mathrm{d}}p$
$=\int_{0}^{1}\left(\gamma^{*}(p)\cdot p+\alpha_{1}^{*}(p)\right)\cdot f_{\mu}(p)\,{\mathrm{d}}p+\int_{0}^{1}(\alpha_{2}^{*}(p)-\gamma^{*}(p)(\kappa_{\nu}(p)-p))\cdot f_{\nu}(p)\,{\mathrm{d}}p$ (due to $\kappa_{\mu}(p)=p$)
$=\int_{0}^{1}\left(\gamma^{*}(p)\cdot p+\alpha_{1}^{*}(p)\right)\cdot f_{\mu}(p)\,{\mathrm{d}}p+\int_{0}^{0.5}(\alpha_{2}^{*}(p)+\gamma^{*}(p)p)\cdot f_{\nu}(p)\,{\mathrm{d}}p+\int_{0.5}^{1}(\alpha_{2}^{*}(p)-\gamma^{*}(p)(1-p))\cdot f_{\nu}(p)\,{\mathrm{d}}p$ (due to $\kappa_{\nu}(p)={\mathbbm{1}\left\{p\geq 0.5\right\}}$)
$=\int_{0}^{1}\left(\gamma^{*}(p)\cdot p+\alpha_{1}^{*}(p)\right)\cdot f_{\mu}(p)\,{\mathrm{d}}p+\int_{0}^{1}\alpha_{2}^{*}(p)f_{\nu}(p)\,{\mathrm{d}}p$
$\quad+\int_{0}^{0.5}\gamma^{*}(p)p\cdot f_{\nu}(p)\,{\mathrm{d}}p-\int_{0.5}^{1}\gamma^{*}(p)(1-p)\cdot f_{\nu}(p)\,{\mathrm{d}}p$
Similar to the analysis of Theorem 3.1, we know that
$\mathcal{L}$ $\geq\sup\limits_{U\text{ is 1-Lipschitz}}\int_{0}^{1}U(p)\cdot(f_{\mu}(p)-f_{\nu}(p))\,{\mathrm{d}}p+\int_{0}^{0.5}U^{\prime}(p)p\cdot f_{\nu}(p)\,{\mathrm{d}}p-\int_{0.5}^{1}U^{\prime}(p)(1-p)\cdot f_{\nu}(p)\,{\mathrm{d}}p$
$\geq\max_{t\in[0,1]}\int_{0}^{1}V_{t}(p)(p)\cdot(f_{\mu}(p)-f_{\nu}(p))\,{\mathrm{d}}p+\int_{0}^{0.5}V^{\prime}_{t}(p)(p)p\cdot f_{\nu}(p)\,{\mathrm{d}}p-\int_{0.5}^{1}V^{\prime}_{t}(p)(p)(1-p)\cdot f_{\nu}(p)\,{\mathrm{d}}p$
$=\max_{t\in[0,1]}4\int_{0}^{\frac{1}{2}}t\cdot f_{\nu}(t)\,{\mathrm{d}}t~{},$
where $V_{t}(p)=|p-t|$. Then by weak duality, we further have
$\overline{\mathrm{REMD}}^{\mathrm{MisC}}\!\left[f_{\mu},f_{\nu},\kappa_{\mu},\kappa_{\nu}\right]\geq 4\int_{0}^{\frac{1}{2}}t\cdot f_{\nu}(t)\,{\mathrm{d}}t~{}.$
We thus finish the proof of Lemma A.1. ∎